### v1

# On completeness of modified wave operators for defocusing NLS

Abstract.
In this manuscript, we study modified scattering for the nonlinear defocusing Schrödinger equation with a critical gauge-invariant nonlinearity of order $1+2/n.$
We address the following question: Given initial data $u_{0}$ in an appropriate weighted Sobolev space, what is the leading term in the asymptotic behavior of the solution $u(t)$ as $t\to\pm\infty$? More precisely, we seek a final state $u^{+}$ as $t\to+\infty$ (in a space of type similar to the space of $u_{0}$) such that
(0.1) $\left\|u(t)-e^{i\Phi^{+}(t)}U(t)u^{+}\right\|_{L^{2}}=0\ \ \mbox{as $t\to+\infty$}.$
This result provides the modified profile of the leading term, where the linear propagator $U(t)$ for the linear Schrödinger equation is perturbed by a modified phase function $\Phi^{+}(t).$
The solution to this problem can be reformulated in terms of the completeness of wave operators. For $n=1$, we obtain a complete answer, even for large initial data. For $n=2$ completeness is established under suitable control of the $L^{\infty}$ norm of the solution $u(t).$
We prove that the $L^{\infty}$ bound of the solution $u(t)$ matches the decay rate of the linear Schrödinger equation only in the case $n=1.$ For $n\geq 2,$ a similar bound holds only for the $L^{2n/(n-1)}$ norm of $u(t).$
In the case where no control on the $L^{\infty}$ norm is available, we establish similar asymptotic behavior by replacing the $L^{2}$ norm in (0.1) with smaller Besov-type spaces.
Keywords: Modified Scattering Nonlinear Schrödinger equation
Subject Class: 35Q55 35P25 35B40
1. Introduction
Consider the Cauchy problem for the defocusing NLS
(1.1) $(i\partial_{t}+\frac{1}{2}\Delta_{x})u=u|u|^{2/n},\ \ t\geq 0,$
$u(0)=u_{0}\in\mathcal{H}^{s,\sigma}=H^{s}(\mathbb{R}^{n})\cap\mathcal{F}(H^{\sigma}(\mathbb{R}^{n})),$
where $u:\mathbb{R}\times\mathbb{R}^{n}\ni(t,x)\mapsto u(t,x)\in\mathbb{C},$ $\partial_{t}=\partial/\partial t,$ $\Delta=\partial_{1}^{2}+\cdots+\partial_{n}^{2},$ $\partial_{j}=\partial/\partial x_{j},$ $H^{s}$ is the classical Sobolev space, see (3.2), and where $\mathcal{F}$ is the Fourier transform introduced in (3.1).
We are interested in the long-range effects of (1.1).
The main questions concerning modified scattering are described below.
  1. Q1)
    Given $u^{+}$ in a space of type $\mathcal{H}^{s,\sigma}$, show that there exists $u_{0}$ and a global solution $u$ to the NLS so that
    (1.2) $\lim_{t\to+\infty}\left\|u(t)-e^{i\Phi^{+}(t)}U(t)u^{+}\right\|_{L^{2}}=0.$
    Here $U(t)=e^{i\frac{t}{2}\Delta}$ is the free propagator, while $\Phi^{+}(t)=\Phi^{+}(t,x)$ is appropriate real -valued function depending on the final state $u^{+}.$ This gives the modified wave operator
    (1.3) $W_{+}(u^{+})=u_{0}=\lim_{t\to+\infty}U(-t)e^{-i\Phi^{+}(t)}u^{+},$
    where the limit is taken in $L^{2}$ sense. Similarly, we can ask if modified wave operator
    (1.4) $W_{-}(u^{-})=u_{0},$
    such that
    (1.5) $\lim_{t\to-\infty}\left\|u(t)-e^{i\Phi^{-}(t)}U(t)u^{-}\right\|_{L^{2}}=0$
    exists?
  2. Q2)
    Another key question is the completeness of wave operators. Given $u_{0}$ in space of type $\mathcal{H}^{\sigma,s},$ suppose that there is a global mild solution $u\in C([0,+\infty);\mathcal{H}^{\sigma,s}).$ One can see for example [34] for the case $n=1$ and [17] for $n=1,2,3$ for the case of initial data in $\mathcal{H}^{0,n/2+}$ and appropriate smallness assumption. Then we are looking for final state $u^{+}$ as $t\to+\infty$ (in a space of type similar to the space of $u_{0}$) such that
    (1.6) $\lim_{t\to+\infty}\left\|u(t)-e^{i\Phi^{+}(t)}U(t)u^{+}\right\|_{L^{2}}=0.$
    In this way we can define
    (1.7) $(W_{+})^{-1}:u_{0}\mapsto u^{+}$
    and in a similar way
    (1.8) $(W_{-})^{-1}:u_{0}\mapsto u^{-}.$
    Further we can introduce the scattering operator
    (1.9) $S=(W_{-})^{-1}W_{+}:u^{+}\mapsto u^{-}.$
The defocussing NLS has conservation quantities: the mass
(1.10) $m(u(t))=\frac{1}{2}\int_{\mathbb{R}^{n}}|u(t,x)|^{2}dx$
and the energy
(1.11) $E(u(t))=\frac{1}{2}\int_{\mathbb{R}^{n}}|\nabla u(t,x)|^{2}dx+\frac{n}{2n+2}\int_{\mathbb{R}^{n}}|u(t,x)|^{2/n+2}dx.$
1.1. Known results
Recall that more general NLS
(1.12) $(i\partial_{t}+\frac{1}{2}\Delta)u=u(t)|u(t)|^{p-1},\ \ t\in[0,+\infty)$
is studied intensively in the recent decades. The critical Strauss exponent
(1.13) $\gamma(n)=\frac{(n+2)+\sqrt{n^{2}+12n+4}}{2n}$
plays crucial role in the proof of existence and completeness of the wave operators. Note that $\gamma(n)>1+2/n$ so this case corresponds to the short range regime.
The case $p>\gamma(n)$ and large initial data in
(1.14) $\mathcal{H}^{1,1}=H^{1}(\mathbb{R}^{n})\cap\mathcal{F}(H^{1}(\mathbb{R}^{n}))$
is studied in [20], [36] and existence and completeness of wave operators is obtained. The nonlinearity with Strauss critical exponent is studied in [7], [8], [31], [32]. The crucial case $p=\gamma(n)$ is studied in [32]. The precise asymptotics for the case $p>\gamma(n)$ is discussed in [27].
Let us mention that most of the known results are treating modified scattering for defocussing critical NLS with $p=1+2/n$ and they need a kind of smallness assumption.
For the special case $n=1$ it is known from the results in [10] that we have the $L^{\infty}$ bound
(1.15) $\|u(t)-t^{-1/2}\alpha(x)e^{ix^{2}/4t-i\nu(x)\log(2t)}\|_{L^{\infty}}\leq C(1+t)^{-1/2-\kappa},\ \kappa\in(0,1/4),$
where $\alpha\in\mathcal{H}^{1,1}$ is complex - valued and $\nu(x)>0.$ The result is obtained by using the inverse scattering method working in dimension $n=1$ and exploiting the reflection coefficient $r(x)$ that is assumed to be in the space
(1.16) $\mathcal{H}^{1,1}\cap\left\{r;\|r\|_{L^{\infty}}<1\right\}.$
This result shows that one can have solution satisfying the decay property
(1.17) $\|u(t,\cdot)\|_{L^{\infty}}\lesssim t^{-1/2},$
when the smallness condition is expressed in terms of explicit bound on the reflection coefficient.
The work [34] is the first case when modified wave operators (1.3), (1.4) for $n=1$ are defined even in the case where the right hand side is replaced by
(1.18) $f(|u(t)|^{2})=\lambda u(t)|u(t)|^{2/n}+\mu u(t)|u(t)|^{p-1},\ \ t\in[0,+\infty),p>1+2/n;$
[…]
(1.19) $\Phi^{\pm}(t,x)=\mp\lambda\log|t|\left|\widehat{u^{\pm}}\left(t^{-1}x\right)\right|^{2}\pm\frac{2\mu}{p-3}|t|^{-(p-3)/2}\left|\widehat{u^{\pm}}\left(t^{-1}x\right)\right|^{p-1},$
$\Phi_{0}^{\pm}(t,x)=\mp\lambda\log|t|\left|\widehat{u^{\pm}}\left(t^{-1}x\right)\right|^{2}.$
The main result in [34] implies (1.3) and (1.4) assuming $u^{\pm}\in\mathcal{H}^{1,3}$ with $\left\|\widehat{u^{\pm}}\right\|_{L^{\infty}}<\varepsilon$. The decay rate in (1.6) is deduced in [34] by verifying the decay estimate
(1.20) $\left\|u(t)-e^{i\Phi^{+}(t)}U(t)u^{+}\right\|_{\mathcal{H}^{1,0}}\lesssim t^{-1+}\ \ \mbox{as $t\to+\infty$}.$
Later on the decay result in [34] has been improved in [5] by using the same assumptions on $u^{+}.$ The cases $n=1,2,3$ are studied in [12], where the Dollard decomposition is used and the existence of wave operators is established. Concerning the Dollard decomposition and modified wave operators we can refer also to [29].
Further, the work [12] treats the Cauchy problem for small and sufficiently regular asymptotic states, i.e. the initial data at $t\to+\infty$ and then the modified scattering is established.
The case of small initial data in the space $\mathcal{H}^{1,1}$ and $p=1+2/n$ is treated in the classical work [16], where the question Q2) is studied and the modified scattering is represented by the following asymptotics as $t\to+\infty$
(1.21) $u(t)=(it)^{-n/2}w\left(\frac{x}{t}\right)e^{ix^{2}/2t+|w(x/t)|^{2}\log t+i\Phi(x/t)}+O(t^{-n/2-\delta}),$
where $w,\Phi$ are $L^{\infty}$ functions. Alternative proof for the case $n=1$ can be found in [24].
The work [17] deals with existence and completeness of the wave operators for $n=1,2,3$ having less regularity of the initial data $u_{0},u^{\pm}\in\mathcal{H}^{0,n/2+}.$ More precisely, if $u_{+}\in\mathcal{H}^{0,\alpha}$ and $\left\|\widehat{u_{+}}\right\|_{{L}^{\infty}}=\varepsilon$, where $\varepsilon$ is sufficiently small and $\frac{n}{2}<\alpha<\min\left(n,2,1+\frac{2}{n}\right)$, then there exists a unique global solution $u$ to the NLS satisfying
(1.22) $\left\|u(t)-\frac{e^{\frac{i|x|^{2}}{2t}}}{(it)^{\frac{n}{2}}}\widehat{u_{+}}\left(\frac{\cdot}{t}\right)\exp\left(-i\lambda\left|\widehat{u_{+}}\left(\frac{\cdot}{t}\right)\right|^{\frac{2}{n}}\log t\right)\right\|_{\mathcal{H}^{\delta,0}}\leq\frac{C}{t^{\frac{\beta-\delta}{2}+\mu}},$
where $\frac{n}{2}<\beta<\alpha$ and $0\leq\delta\leq\beta,\mu>0$.
Further the operator $W_{-}^{-1}$ is constructed in [17], so that
$W_{-}^{-1}:u_{0}\in\mathcal{H}^{0,\beta}\mapsto u_{-}\in\mathcal{H}^{0,\delta}$
where $\frac{n}{2}<\delta<\beta<\min\left(2,1+\frac{2}{n}\right)$. Therefore we have the modified scattering operator
$S_{+}=W_{-}^{-1}W_{+}:\mathcal{H}^{0,\alpha}\mapsto\mathcal{H}^{0,\delta},$
where $\frac{n}{2}<\delta<\alpha<\min\left(n,2,1+\frac{2}{n}\right)$, provided that the norm $\left\|u_{+}\right\|_{\mathcal{H}^{0,\alpha}}$ is sufficiently small.
The work [28] gives another interesting result for 1d long range scattering of NLS. The authors make the ansatz
$u(t,x)=t^{-1/2}\mathrm{e}^{\mathrm{i}x^{2}/4t}V(t,y),\quad s=t,\quad y=x/t$
Then the NLS problem is transformed into
(1.23) $\mathrm{i}\partial_{s}V-\beta s^{-1}|V|^{2}V+s^{-2}\partial_{y}^{2}V=0$
and then $V$ is represented as $V=ae^{i\phi}.$ Then the authors prove the existence and completeness of wave operators under smallness assumption of the smooth initial data.
In [30] modified 1d scattering is established under smallness assumptions in $\mathcal{H}^{1,1}.$
The asymptotic expansion with respect to the initial small data size is given in [6].
On the other hand, if initial data are in $H^{1}$ only, then the following estimate is known (see [35], [9]) for the 1d case
(1.24) $\|u\|_{L^{6}_{(1,+\infty)}L^{6}_{x}}^{2}+\left\|\nabla_{x}(|u|^{2})\right\|_{L^{2}_{(1,+\infty)}L^{2}_{x}}\lesssim\sup_{t\geq 1}\|u(t)\|_{L^{2}}\|u(t)\|_{\dot{H}^{1/2}}.$
Using smallness assumption
(1.25) $\left\|u_{0}\right\|_{\mathcal{H}^{0,1}}\leq\varepsilon<<1,$
Ifrim and Tataru in [22] have obtained the smallness of $L^{\infty}$ norm and the expected decay $t^{-1/2}.$ Note that the smallness assumption (1.25) can be weaken to the smallness assumption of the $\mathcal{H}^{0,1/2+}$ and this is exactly the case treated in [17]. The decay of $L^{\infty}$ norm is studied in [19] under the assumption that $p\geq 1+4/n.$ In the case $n=1$ the small weighted Sobolev norms on initial data (actually exponential decay in $x$ is assumed) the modified scattering is deduced in [15].
In the case of more general cubic nonlinearities and $n=1$ the existence of wave operators are established in [18] by requiring $u_{0}\in\mathcal{H}^{2,2}.$
1.2. Purpose of the work
The above short overview on existing results shows that most of the results require smallness of the initial data
Our goal is to solve Q2) without smallness on the initial data at least for $n=1.$
Remark 1.1.
Let us mention that in the case $n=1$ the question Q1) is treated in the preprint [25].
2. Main results
Theorem 2.1.
Assume $n\leq 2$ and let $u(t)$ be a global solution to (1.1) satisfying
(2.1) $\|u(t,\cdot)\|_{L^{\infty}}\lesssim t^{-n/2},\ \ t\geq 1$
with $u(0)=u_{0}\in\mathcal{F}(H^{1}).$ Then there exists a unique $u^{+}\in L^{2}$ so that
(2.2) $\lim_{t\to+\infty}\|u(t)-e^{-i\Phi_{t}(u)}U(t)u^{+}\|_{L^{2}}=0,$
where
(2.3) $\Phi_{t}(u)(x)=\int_{1}^{t}|u(\tau,\tau x/t)|^{2/n}d\tau.$
Remark 2.1.
From the results, treating 1d case in [26], we know that the modified wave operators
(2.4) $W_{+}:u^{+}\in\mathcal{F}(H^{1+\varepsilon})\mapsto u_{0}=\lim_{t\to+\infty}U(-t)e^{-i\Phi^{+}(t)}u^{+}\in\mathcal{F}(H^{1}),$
$W_{-}:u^{-}\in\mathcal{F}(H^{1+\varepsilon})\mapsto u_{0}=\lim_{t\to+\infty}U(-t)e^{-i\Phi^{-}(t)}u^{-}\in\mathcal{F}(H^{1})$
are well defined and any $u_{0}\in\mathrm{Ran}(W_{-})$ satisfies the decay assumption (2.1). From this fact and from Theorem 2.1 we can conclude that
  • the modified wave operators are complete,
  • the scattering operator $S=(W_{+})^{-1}W_{-}$ is well - defined.
Theorem 2.2.
Under the assumptions of Theorem 2.1 if in addition $u_{0}\in H^{1},$ then $u^{+}\in H^{1}$ and for any $\varphi\in H^{1}$ we have
(2.5) $\lim_{t\to+\infty}\left(\nabla\left(U(-t)e^{i\Phi_{t}(u)}u(t)-u^{+}\right)|\nabla\varphi\right)=0,$
where $(\cdot|\cdot)$ is the inner product in $L^{2}$ defined by (3.4).
Remark $2.$
The limit in (2.5) and (2.2) show that the weak limit in $H^{1}$ of
$U(-t)e^{i\Phi_{t}(u)}u(t)-u^{+}$
is $0.$
Let us give an outline of the proofs of these Theorems. We use the Dollard decomposition (see [12], [29]) of the Schrödinger group $U(t)=e^{i\frac{t}{2}\Delta}$ represented by
(2.6) $(U(t)f)(x)=\int_{\mathbb{R}^{n}}K_{t}(x-y)f(y)dy,$
where
(2.7) $K_{t}(x)=(2\pi it)^{-n/2}e^{ix^{2}/(2t)},t\neq 0.$
The Dollard decomposition has the form
(2.8) $U(t)=M(t)D(t)\mathcal{F}M(t),$
where
(2.9) $M(t)(f)(x)=e^{ix^{2}/(2t)}f(x),\ \ M(t)^{-1}=M(-t),$
$D(t)\,(f)(x)=(it)^{-n/2}f\left(\frac{x}{t}\right),\ D(t)^{-1}\,(f)(x)=(it)^{n/2}f\left(xt\right).$
The proofs then rely on the decay assumption and the pseudoconformal energy. To establish estimates for the pseudoconformal energy, we apply the pseudoconformal transformation.
(2.10) $u(t,x)=(1+t)^{-n/2}v\left(\frac{t}{1+t},\frac{x}{1+t}\right)e^{\frac{ix^{2}}{2(1+t)}},$
well - defined for $t\geq 0,x\in\mathbb{R}^{n}.$ Then $v\in C([0,1);\mathcal{H}^{1,1})$ is a solution to the Cauchy problem
(2.11) $(i\partial_{t}+\frac{1}{2}\Delta_{x})v=\frac{v|v|^{2/n}}{1-t},\ \ t\in[0,1);$
$v(0)=u_{0}\in\mathcal{H}^{1,1}.$
Next, our goal is to prove some $L^{q}$ decay estimates closely connected with the key assumption (2.1) in Theorem 2.1.
Theorem 2.3.
If $n=1$ and $u_{0}\in\mathcal{H}^{1,1},$ then the solution to (1.1) satisfies
(2.12) $\|u(t)\|_{L^{\infty}}\leq Ct^{-1/2},t\geq 1.$
For $n\geq 2$ we can control the $L^{q}$ norm of the solution, where $q$ is limited by an index associated with Caffarelli - Kohn - Nirenberg inequality.
Theorem 2.4.
If $n\geq 2$, $u_{0}\in\mathcal{H}^{1,1},$ and $q\in[2,2n/(n-1)],$ then the solution to (1.1) satisfies
(2.13) $\|u(t)\|_{L^{q}}\leq Ct^{-n/2+n/q},t\geq 1.$
We have another estimate provided initial data are radial.
Theorem 2.5.
If $n\geq 2$ and $u_{0}\in\mathcal{H}^{1,1}_{rad},$ then the solution to (1.1) satisfies
(2.14) $\||x|^{n/2}u(t)\|_{L^{\infty}}\leq C,t\geq 1.$
Again we give a brief idea of the proofs of these $L^{q}$ bounds.
We use pseudoconformal transformation (2.10). Then the results of Theorems 2.3, 2.4, 2.2 follow from the following.
Theorem 2.6.
If $n=1$ and $u_{0}\in\mathcal{H}^{1,1},$ then the solution to (2.11) satisfies
(2.15) $\|v(t)\|_{L^{\infty}}\leq C,t\in[0,1).$
Theorem 2.7.
If $n\geq 2$, $u_{0}\in\mathcal{H}^{1,1}$ and $q\in[2,2n/(n-1)],$ then the solution to (2.11) satisfies
(2.16) $\|v(t)\|_{L^{q}}\leq C,t\in[0,1).$
Theorem 2.8.
If $n\geq 2$ and $u_{0}\in\mathcal{H}^{1,1}_{rad},$ then the solution to (2.11) satisfies
(2.17) $\||x|^{n/2}v(t)\|_{L^{\infty}}\leq C,t\in[0,1).$
Then the $L^{q}$ bounds are obtained by using the energy estimate for $v$ (see Lemma 6.2)
(2.18) $\int_{\mathbb{R}^{n}}|\nabla_{x}v(t,x)|^{2}dx\leq C(1-t)^{-1}$
and the virial estimate;
(2.19) $\int_{\mathbb{R}^{n}}|x|^{2}|v(t,x)|^{2}dx\leq C(1-t).$
A natural interpretation of this estimate is that the solution is spreading out (dispersing) with a very strong rate.
The crucial estimate (2.19) is obtained via appropriate half - pseudoconformal transformation
(2.20) $v(t,x)=e^{-\frac{ix^{2}}{4(1-t)}}(1-t)^{-n/4}Y\left(t,\frac{x}{\sqrt{1-t}}\right).$
Then $Y(t,y)$ satisfies a non - linear Schrödinger equation involving inverted harmonic oscillator. We use a combination between virial identity for this equation and energy estimate involving the vector field
(2.21) $\partial_{y}-i\frac{y}{2}$
that allows us to obtain
(2.22) $\int_{\mathbb{R}^{n}}|y|^{2}|Y(t,y)|^{2}dy\leq C$
and then we arrive at (2.19).
In addition, we turn to the case where no $L^{\infty}$ control is guaranteed. In the case $n=2$ we have the following result that is a generalization of a result of Hirata in [21]. We use the standard notation of homogeneous Besov spaces $\dot{B}^{s}_{p,q}$ (see [2] for definition and properties).
Theorem 2.9.
For any $u_{0}\in\mathcal{F}(H^{1})$ there exists a unique $u_{+}\in L^{2},$ so that
(2.23) $\lim_{t\to+\infty}\left\|U(-t)\exp\left(i\Phi_{t}(u)\right)u(t)-u_{+}\right\|_{\mathcal{F}(\dot{B}^{-1}_{2,\infty})}=0,$
where $\Phi_{t}$ is defined in (2.3).
Turning to the case $n\geq 3,$ we consider the problem
(2.24) $(i\partial_{t}+\frac{1}{2}\Delta_{x})u=u((1+|u|^{2})^{1/n}-1),\ \ t\geq 0;$
$u(0)=u_{0}.$
The nonlinearity is associated with the functions.
(2.25) $f(z)=(\langle z\rangle^{2/n}-1)z,$
$V(z)=\frac{n}{n+1}(\langle z\rangle^{2(n+1)/n}-1)-|z|^{2},$
$W(z)=(n+2)V(z)-n\overline{z}f(z)$
so that $f(z)=\partial_{\overline{z}}V$ and $f(0)=V(0)=W(0)=0.$
Theorem 2.10.
For any $u_{0}\in\mathcal{F}(H^{1})$ there exists a unique $u_{+}\in L^{2},$ so that
(2.26) $\lim_{t\to+\infty}\left\|U(-t)\exp\left(i\Phi_{t}(u)\right)u(t)-u_{+}\right\|_{\mathcal{F}(\dot{B}^{-n/2}_{2,\infty})}=0,$
where $u(t)$ is the solution to (2.24).
3. Notations
The Fourier transform is defined by
(3.1) $\mathcal{F}\varphi(\xi)=(2\pi)^{-n/2}\int_{\mathbb{R}^{n}}e^{-ix\cdot\xi}\varphi(x)dx$
for $\varphi\in\mathcal{S},$ rapidly decreasing $C^{\infty}$ functions. Here and below $x\cdot\xi$ means the scalar product of $x,\xi\in\mathbb{R}^{n}$ in $\mathbb{R}^{n}.$ Often, we shall write $x^{2}$ in the place of $x\cdot x$ for $x\in\mathbb{R}^{n}.$
We use the notation $H^{s}$ for the classical Sobolev space
(3.2) $H^{s}=H^{s}(\mathbb{R}^{n})=(1-\Delta)^{-s/2}L^{2}(\mathbb{R}^{n})$
$=\mathcal{F}^{-1}(1+|\xi|^{2})^{-s/2}\mathcal{F}L^{2}(\mathbb{R}^{n}).$
For any $z=(z_{1},\cdots,z_{k})\in\mathbb{C}^{k}$ we use the notation
(3.3) $\langle z\rangle=(1+|z|^{2})^{1/2},|z|^{2}=\sum_{j=1}^{k}|z_{j}|^{2}.$
In the Hilbert space $L^{2}(\mathbb{R}^{n};\mathbb{C}),$ which consists of square-integrable complex-valued functions on $\mathbb{R}^{n},$ the inner product is defined as:
(3.4) $\left.(f\right|g)=\int_{\mathbb{R}^{n}}f(x)\overline{g(x)}dx.$
The inequality $X\lesssim Y$ is used as a shorthand notation for inequalities that involve implicit constants. Specifically, $X\lesssim Y$ means $X\leq CY$ for some constant $C.$
4. Preliminary calculations
We use the Dollard decomposition (2.8), involving the operators $M(t),D(t)$ defined in (2.9).
We have the relations
(4.1) $U(-t)=M(t)^{-1}\mathcal{F}^{-1}D(t)^{-1}M(t)^{-1},$
$\mathcal{F}M(t)U(-t)=D(t)^{-1}M(t)^{-1}.$
Note that multiplication by a function commutes with $M(t)$
(4.2) $M(t)(gf)=gM(t)(f).$
Moreover,
(4.3) $(D(t)^{-1}\ e^{i\Phi_{t}}f)(x)=(it)^{n/2}e^{i\Phi_{t}(xt)}f\left(xt\right)$
$=e^{i\tilde{\Phi}_{t}(x)}(D(t)^{-1}f)(x),$
where
(4.4) $\tilde{\Phi}_{t}(x)=\int_{1}^{t}|u(\tau,\tau x)|^{2/n}d\tau.$
Then we set
(4.5) $v(t)=M(t)U(-t)e^{i\Phi_{t}(u)}u(t).$
It is easy to deduce the following.
Lemma 4.1.
For any $g\in\mathcal{S}$ we have
(4.6) $\mathcal{F}M(t)U(-t)(gu(t))=\tilde{g}\mathcal{F}M(t)U(-t)u(t),\ \ \ \tilde{g}(x)=g(xt).$
Proof.
It is sufficient to apply (4.1) that implies
(4.7) $\mathcal{F}M(t)U(-t)=D(t)^{-1}M(t)^{-1}.$
Then we can quote the following variant of (4.3).
(4.8) $D(t)^{-1}(gf)(x)=\tilde{g}D(t)^{-1}(f)(x)$
and complete the proof. ∎
We can apply the above Lemma and from (4.5) we can write
(4.9) $v(t)=\mathcal{F}^{-1}\mathcal{F}M(t)U(-t)e^{i\Phi_{t}(u)}u(t)=\mathcal{F}^{-1}e^{i\tilde{\Phi}_{t}}\mathcal{F}M(t)U(-t)u(t)$
due to (4.1).
In this way, we get
(4.10) $i\partial_{t}v(t)=\mathcal{F}^{-1}i\partial_{t}\left(e^{i\tilde{\Phi}_{t}}\mathcal{F}M(t)U(-t)u(t)\right)$
$=-\mathcal{F}^{-1}\partial_{t}\tilde{\Phi}_{t}e^{i\tilde{\Phi}_{t}}\mathcal{F}M(t)U(-t)u(t)$
$+i\mathcal{F}^{-1}e^{i\tilde{\Phi}_{t}}\mathcal{F}\partial_{t}\left(M(t)U(-t)u(t)\right).$
On the other hand, we have the integral representation
(4.11) $u(t)=U(t)u_{0}-i\int_{0}^{t}U(t-\tau)\ \ u(\tau)|u(\tau)|^{2/n}\ \ d\tau.$
Hence
(4.12) $i\partial_{t}\left(U(-t)u(t)\right)=U(-t)u(t)|u(t)|^{2/n}.$
Further we have the simple relation
(4.13) $\partial_{t}(M(t)w(t))=M(t)\left(\partial_{t}w(t)-i\frac{x^{2}}{2t^{2}}w(t)\right)$
hence
(4.14) $i\partial_{t}\left(M(t)U(-t)u(t)\right)=M(t)U(-t)\left(u(t)|u(t)|^{2/n}\right)+\frac{x^{2}}{2t^{2}}M(t)U(-t)u(t).$
Now we turn back to (4.10) and we can write
(4.15) $i\partial_{t}v(t)=-\mathcal{F}^{-1}e^{i\tilde{\Phi}_{t}}|u(t,xt)|^{2/n}\mathcal{F}M(t)U(-t)u(t)$
$+\mathcal{F}^{-1}e^{i\tilde{\Phi}_{t}}\mathcal{F}M(t)U(-t)\left(u(t)|u(t,x)|^{2/n}\right)$
$+\mathcal{F}^{-1}e^{i\tilde{\Phi}_{t}}\mathcal{F}\frac{x^{2}}{2t^{2}}M(t)U(-t)u(t)$
$=-\mathcal{F}^{-1}e^{i\tilde{\Phi}_{t}}\mathcal{F}M(t)U(-t)\left(|u(t,xt)|^{2/n}u(t)\right)$
$+\mathcal{F}^{-1}e^{i\tilde{\Phi}_{t}}\mathcal{F}M(t)U(-t)\left(u(t)|u(t,x)|^{2/n}\right)$
$+\mathcal{F}^{-1}e^{i\tilde{\Phi}_{t}}\mathcal{F}\frac{x^{2}}{2t^{2}}M(t)U(-t)u(t)$
$=-\frac{1}{2t^{2}}\mathcal{F}^{-1}e^{i\tilde{\Phi}_{t}}\Delta\mathcal{F}M(t)U(-t)u(t).$
Thus we obtain
Corollary 4.1.
(4.16) $i\frac{d}{dt}v(t)=i\frac{d}{dt}\left(M(t)U(-t)\exp\left(i\int_{0}^{t}|u(\tau,\tau x/t)|^{p-1}d\tau\right)u(t)\right)$
$=\mathcal{F}^{-1}\exp\left(i\int_{0}^{t}|u(\tau,\tau x)|^{p-1}d\tau\right)\frac{(-\Delta)}{2t^{2}}D(t)^{-1}M(-t)u(t)$
$=\mathcal{F}^{-1}\exp\left(i\int_{0}^{t}|u(\tau,\tau x)|^{p-1}d\tau\right)\frac{(-\Delta)}{2t^{2}}\mathcal{F}M(t)U(-t)u(t).$
5. $L^{\infty}$ bound implies the existence of modified wave operator
Proof of Theorem 2.1 .
We start with the plan of the proof. It consists of the following steps.
  1. Step I)
    We have to prove that
    (5.1) $v(t)=M(t)U(-t)e^{i\Phi_{t}(u)}u(t)$
    has $L^{2}$ limit as $t\to+\infty.$ Indeed, the $L^{2}$ limit in (2.2) is valid if
    $U(-t)e^{i\Phi_{t}(u)}u(t)$
    is $L^{2}$ Cauchy sequence as $t\to+\infty.$ Taking into account the fact that
    (5.2) $\lim_{t\to+\infty}\|M(t)g-g\|_{L^{2}}=0,\ \ \forall g\in L^{2}$
    due to Lebesgue convergence theorem, we see that we have to check that $v(t)$ is a Cauchy sequence. Therefore in this step we shall verify this property assuming $n\leq 2$ we have the $L^{\infty}$ control (2.1) as well as we have the uniform estimates
    (5.3) $\|J(t+1)u(t)\|_{L^{2}}\lesssim(1+t)^{1/2}\|u(0)\|_{\mathcal{H}^{0,1}}.$
    (5.4) $\int_{0}^{t}(1+\tau)^{-2}\|J(1+\tau)u(\tau)\|_{L^{2}}^{2}d\tau\lesssim\|u(0)\|^{2}_{\mathcal{H}^{0,1}}.$
    Here and below $J(1+t)$ is the generator of the pseudoconformal group
    (5.5) $J(1+t)=x+i(1+t)\nabla_{x};$
  2. Step II)
    We prove (5.3) and (5.4) by using pseudoconformal transform and applying simple energy estimate.
Now we turn to the details of the proof of Theorem 2.1.
Note that $v(t)$ satisfies the relations (see Lemma 4.1)
(5.6) $\mathcal{F}(v(t))=\mathcal{F}M(t)U(-t)e^{i\Phi_{t}(u)}u(t)$
$=e^{i\tilde{\Phi}_{t}}\mathcal{F}M(t)U(-t)u(t)=e^{i\tilde{\Phi}_{t}}\mathcal{F}M(t)U(-t)u(t).$
The conservation of mass guarantees that it is sufficient to prove that
(5.7) $\frac{\|v(t)-v(s)\|_{L^{2}}^{2}}{2}=\|v(s)\|_{L^{2}}^{2}-\left.\mathrm{Re}(v(t)\right||v(s))$
is small if $t>s>M$ and if $M$ is large.
We use Corollary 4.1 and setting
(5.8) $\mathbb{W}(t)=M(t)U(-t)u(t),$
we can write
(5.9) $\|v(s)\|_{L^{2}}^{2}-\mathrm{Re}\left(v(t)\big|v(s)\right)=-\mathrm{Re}\left(v(t)-v(s)|v(s)\right)$
$=-\mathrm{Re}\int_{s}^{t}\left(v^{\prime}(\tau)|v(s)\right)d\tau$
$=\left.\mathrm{Im}\int_{s}^{t}\left(\mathcal{F}^{-1}\exp\left(i\tilde{\Phi}_{\tau}\right)\frac{(-\Delta)}{2\tau^{2}}\mathcal{F}\mathbb{W}(\tau)\right|M(s)U(-s)e^{i\Phi_{s}}u(s)\right)d\tau$
$=\left.\mathrm{Im}\int_{s}^{t}\left(\exp\left(i\tilde{\Phi}_{\tau}\right)\frac{(-\Delta)}{2\tau^{2}}\mathcal{F}\mathbb{W}(\tau)\right|e^{i\tilde{\Phi}_{s}}\mathcal{F}\mathbb{W}(s)\right)d\tau$
$=\left.\mathrm{Im}\int_{s}^{t}\left(\exp\left(i\tilde{\Phi}_{\tau}-i\tilde{\Phi}_{s}\right)\frac{(-\Delta)}{2\tau^{2}}\mathcal{F}\mathbb{W}(\tau)\right|\mathcal{F}\mathbb{W}(s)\right)d\tau$
$=\left.\mathrm{Im}\int_{s}^{t}\left(\exp\left(i\tilde{\Phi}_{\tau}-i\tilde{\Phi}_{s}\right)\frac{\nabla_{x}}{2\tau^{2}}\mathcal{F}\mathbb{W}(\tau)\right|\nabla_{x}\mathcal{F}\mathbb{W}(s)\right)d\tau$
$+\left.\mathrm{Im}\int_{s}^{t}\left(\nabla_{x}\left[\exp\left(i\tilde{\Phi}_{\tau}-i\tilde{\Phi}_{s}\right)\right]\frac{\nabla_{x}}{2\tau^{2}}\mathcal{F}\mathbb{W}(\tau)\right|\mathcal{F}\mathbb{W}(s)\right)d\tau.$
Setting
(5.10) $I(t,s)=\left.\mathrm{Im}\int_{s}^{t}\left(\exp\left(i\tilde{\Phi}_{\tau}-i\tilde{\Phi}_{s}\right)\frac{\nabla_{x}}{2\tau^{2}}\mathcal{F}\mathbb{W}(\tau)\right|\nabla_{x}\mathcal{F}\mathbb{W}(s)\right)d\tau,$
$II(t,s)=\left.\mathrm{Im}\int_{s}^{t}\left(\nabla_{x}\left[\exp\left(i\tilde{\Phi}_{\tau}-i\tilde{\Phi}_{s}\right)\right]\frac{\nabla_{x}}{2\tau^{2}}\mathcal{F}\mathbb{W}(\tau)\right|\mathcal{F}\mathbb{W}(s)\right)d\tau,$
where here and below we use the definition of $\mathbb{W}$ from (5.8). We can estimate the first term as follows
(5.11) $|I(t,s)|\lesssim\int_{s}^{t}\left\|\frac{(\nabla_{x})}{2\tau^{2}}\mathcal{F}\mathbb{W}(\tau)\right\|_{L^{2}}\left\|\nabla_{x}\mathcal{F}\mathbb{W}(s)\right\|_{L^{2}}d\tau.$
Note that
(5.12) $\left\|\nabla_{x}\mathcal{F}\mathbb{W}(\tau)\right\|_{L^{2}}=\left\|\nabla_{x}\mathcal{F}M(\tau)U(-\tau)u(\tau)\right\|_{L^{2}}$
$=\left\|xM(\tau)U(-\tau)u(\tau)\right\|_{L^{2}}=\left\|U(\tau)xU(-\tau)u(\tau)\right\|_{L^{2}}=\left\|J(\tau)u(\tau)\right\|_{L^{2}},$
where
(5.13) $J(t)=U(t)xU(-t)=x+it\nabla_{x}$
and
(5.14) $\left\|J(\tau)u(\tau)\right\|_{L^{2}}\leq\left\|J(1+\tau)u(\tau)\right\|_{L^{2}}+\left\|\nabla u(\tau)\right\|_{L^{2}}$
$\lesssim\tau^{1/2}+O(1),\ \ \tau\to\infty$
due to uniform bounds of Proposition 6.1. Then we find
(5.15) $|I(t,s)|\lesssim\int_{s}^{t}\tau^{-2}\left\|J(\tau)u(\tau)\right\|_{L^{2}}\left\|Ju(s)\right\|_{L^{2}}d\tau$
$\lesssim s^{1/2}\int_{s}^{t}\tau^{-2}\left\|J(\tau)u(\tau)\right\|_{L^{2}}d\tau$
$\lesssim s^{1/2}\left(\int_{s}^{t}\frac{d\tau}{\tau^{2}}\right)^{1/2}\left(\int_{s}^{t}\left\|J(\tau)u(\tau)\right\|_{L^{2}}^{2}\frac{d\tau}{\tau^{2}}\right)^{1/2}=o(1)$
as $s\to+\infty.$ To estimate the second term we proceed as follows
(5.16) $|II(t,s)|\lesssim\int_{s}^{t}\left\|\nabla_{x}\left(i\tilde{\Phi}_{\tau}-i\tilde{\Phi}_{s}\right)\right\|_{L^{2}}\left\|\frac{(\nabla_{x})}{2\tau^{2}}\mathcal{F}\mathbb{W}(\tau)\right\|_{L^{2}}\|\mathcal{F}\mathbb{W}(s)\|_{L^{\infty}}d\tau.$
On the other hand, we use (4.1) and deduce
(5.17) $\|\mathcal{F}\mathbb{W}(s)\|_{L^{\infty}}=\|\mathcal{F}M(s)U(-s)u(s)\|_{L^{\infty}}=\|D(-s)M(-s)u(s)\|_{L^{\infty}}$
$=\|D(-s)u(s)\|_{L^{\infty}}=s^{n/2}\|u(s)\|_{L^{\infty}}=O(1).$
Further, we can write
(5.18) $\nabla_{x}\left(u(\sigma,\sigma x)e^{ix^{2}\sigma/2}\right)$
$=e^{ix^{2}\sigma/2}\left[\sigma(\nabla_{y}u)(\sigma,\sigma x)+i\sigma xu(\sigma,\sigma x)\right]$
$=ie^{ix^{2}\sigma/2}(J(\sigma)u(\sigma))(\sigma x),$
so we get
(5.19) $\left|\nabla_{x}\left(i\tilde{\Phi}_{\tau}-i\tilde{\Phi}_{s}\right)\right|\lesssim\int_{s}^{\tau}\left|\nabla|u(\sigma,\sigma x)|^{2/n}\right|d\sigma$
$=\int_{s}^{\tau}\left|\nabla_{x}|e^{-ix^{2}\sigma/4}u(\sigma,\sigma x)|^{2/n}\right|d\sigma$
$\lesssim\int_{s}^{\tau}\left|\nabla_{x}|e^{-ix^{2}\sigma/4}u(\sigma,\sigma x)|^{2/n}\right|d\sigma$
$\lesssim\int_{s}^{\tau}|(J(\sigma)u(\sigma))(\sigma x)||u(\sigma,\sigma x)|^{2/n-1}d\sigma$
$\lesssim\int_{s}^{\tau}|(J(\sigma)u(\sigma))(\sigma x)|\sigma^{-1+n/2}d\sigma.$
Taking the $L^{2}$ norm we find
(5.20) $\left\|\nabla_{x}\left(i\tilde{\Phi}_{\tau}-i\tilde{\Phi}_{s}\right)\right\|_{L^{2}}$
$\lesssim\int_{s}^{\tau}\|(J(\sigma)u(\sigma))(\sigma\cdot)\|_{L^{2}}\ \sigma^{-1+n/2}d\sigma$
$=\int_{s}^{\tau}\|J(\sigma)u(\sigma)\|_{L^{2}}\frac{d\sigma}{\sigma}.$
Therefore, using (5.14), we arrive at
(5.21) $|II(t,s)|\lesssim\int_{s}^{t}\tau^{-2}\int_{s}^{\tau}\|J(\sigma)u(\sigma)\|_{L^{2}}\frac{d\sigma}{\sigma}\left\|J(\tau)u(\tau)\right\|_{L^{2}}d\tau$
$\lesssim\left(\int_{s}^{t}\tau^{-2}\left(\int_{s}^{\tau}\|J(\sigma)u(\sigma)\|_{L^{2}}\frac{d\sigma}{\sigma}\right)^{2}d\tau\right)^{1/2}\left(\int_{s}^{t}\tau^{-2}\left\|J(\tau)u(\tau)\right\|_{L^{2}}^{2}d\tau\right)^{1/2}.$
Now recall a minor modification of the classical Hardy inequality.
Lemma 5.1.
If $F(s)\in C(0,+\infty)\cap L^{2}(0,+\infty),$ then
(5.22) $\int_{s}^{t}\tau^{-2}\left(\int_{s}^{\tau}F(\sigma)d\sigma\right)^{2}d\tau\leq 4\int_{s}^{t}F(\tau)^{2}d\tau$
for $0<s<t<+\infty.$
Proof.
Set $f(\tau)=F(s+\tau).$ Then the inequality (5.22) is transformed into
(5.23) $\int_{0}^{t-s}(s+\tau)^{-2}\left(\int_{0}^{\tau}f(\sigma)d\sigma\right)^{2}d\tau\leq 4\int_{0}^{t-s}f(\tau)^{2}d\tau$
Since
(5.24) $(s+\tau)^{-1}<\tau^{-1},$
we see that the classical Hardy inequality (see Chapter IX, section 327, inequality (9.8.2) in [14])
(5.25) $\int_{0}^{+\infty}\tau^{-2}\left(\int_{0}^{\tau}f(\sigma)d\sigma\right)^{2}d\tau\leq 4\int_{0}^{+\infty}f(\tau)^{2}d\tau$
implies (5.23). This completes the proof. ∎
We apply (5.22) with $F(\sigma)=\sigma^{-1}\|J(\sigma)u(\sigma)\|_{L^{2}}$ and via (5.14) we arrive at
(5.26) $|II(t,s)|\lesssim\int_{s}^{t}\tau^{-2}\left\|J(\tau)u(\tau)\right\|_{L^{2}}^{2}d\tau=o(1).$
6. Proof of (5.3) and (5.4) via pseudoconformal transformation
6.1. Main Proposition.
The key point in this section is the following.
Proposition 6.1.
The solution $u$ to (1.1) satisfies
  1. (1)
    the conservation of the mass
    (6.1) $\|u(t)\|_{L^{2}}^{2}=\|u_{0}\|_{L^{2}}^{2},$
  2. (2)
    the conservation of the pseudoconformal energy
    (6.2) $(1+t)^{-1}\|J(1+t)u(t)\|_{L^{2}}^{2}+\frac{2n}{n+1}(1+t)\|u(t)\|_{L^{2+2/n}}^{2+2/n}$
    $+\int_{0}^{t}(1+s)^{-2}\|Ju(s)\|_{L^{2}}^{2}ds$
    $=\|J(1)u(0)\|_{L^{2}}^{2}+\frac{2n}{n+1}\|u(0)\|_{L^{2+2/n}}^{2+2/n},$
    where
    (6.3) $J(t+1)=x+i(1+t)\nabla.$
Remark 6.1.
Since we use pseudoconformal transformation, we follow [11] and the method of the proof in this work is based on the approach in [33].
6.2. Pseudoconformal transformation
The first known application of the pseudoconformal transformation to the NLS dates back to the 1980s, (see [13], [37], [38], [8]). We use the following pseudoconformal transformation $(t,x,u)\mapsto(\tau,\xi,v)$ defined by
(6.4) $t=\frac{\tau}{1-\tau},\ x=\frac{\xi}{1-\tau},$
$u(t,x)=(1+t)^{-n/2}v\left(\frac{t}{1+t},\frac{x}{1+t}\right)e^{\frac{ix^{2}}{2(1+t)}}$
with inverse transformation given by
(6.5) $\tau=\frac{t}{1+t},\ \xi=\frac{x}{1+t},$
$v(\tau,\xi)=u\left(\frac{\tau}{1-\tau},\frac{\xi}{1-\tau}\right)(1-\tau)^{-n/2}e^{-\frac{i\xi^{2}}{2(1-\tau)}}.$
If $u$ is a solution to (1.1), then we can apply the pseudoconformal transformation and we can obtain
(6.6) $(i\partial_{\tau}+\frac{1}{2}\Delta_{\xi})v=\frac{v(\tau)|v(\tau)|^{2/n}}{1-\tau},\ \ \tau\in[0,1),$
$v(0)=v_{0},$
where
(6.7) $v_{0}(x)=u_{0}(x)e^{-ix^{2}/2}.$
Lemma 6.1.
We have the relations
(6.8) $\|u(t)\|_{L^{2}}=\|v(\tau)\|_{L^{2}},\$
$\left\|u\left(t\right)\right\|^{2(n+1)/n}_{L^{2(n+1)/n}}=(1+t)^{-1}\left\|v\left(\tau\right)\right\|^{2(n+1)/n}_{L^{2(n+1)/n}},$
$\|\xi v(\tau)\|_{L^{2}}=(1+t)^{-1}\|xu(t)\|_{L^{2}},$
$\|\nabla_{\xi}v(\tau)\|_{L^{2}}=\|J(t+1)u(t)\|_{L^{2}},$
where
(6.9) $\tau=\frac{t}{1+t},\ \ J(t+1)=x+i(1+t)\nabla.$
Proof.
From (6.4) we have
(6.10) $|u(t,x)|=\left|v\left(\frac{t}{1+t},\frac{x}{1+t}\right)\right|(1+t)^{-n/2}=\left|v\left(\tau,\frac{x}{1+t}\right)\right|(1+t)^{-n/2},$
and we find
(6.11) $\int_{\mathbb{R}^{n}}|u(t,x)|^{2}dx=\int_{\mathbb{R}^{n}}(1+t)^{-n}\left|v\left(\tau,\frac{x}{1+t}\right)\right|^{2}dx=\left\|v\left(\tau\right)\right\|^{2}_{L^{2}}$
and
(6.12) $\int_{\mathbb{R}^{n}}|u(t,x)|^{q}dx=\int_{\mathbb{R}^{n}}(1+t)^{-nq/2}\left|v\left(\tau,\frac{x}{1+t}\right)\right|^{q}dx$
$=(1+t)^{n-nq/2}\left\|v\left(\tau\right)\right\|^{q}_{L^{q}}.$
Taking $q=2+2/n=2(n+1)/n,$ we find
(6.13) $\left\|u\left(t\right)\right\|^{2(n+1)/n}_{L^{2(n+1)/n}}=(1+t)^{-1}\left\|v\left(\tau\right)\right\|^{2(n+1)/n}_{L^{2(n+1)/n}}.$
In a similar way, we have
(6.14) $\int_{\mathbb{R}^{n}}x^{2}|u(t,x)|^{2}dx=\int_{\mathbb{R}^{n}}(1+t)^{-n+2}\left|\frac{x}{1+t}v\left(\tau,\frac{x}{1+t}\right)\right|^{2}dx$
$=(1+t)^{2}\left\||\xi|v\left(\tau\right)\right\|^{2}_{L^{2}}.$
Finally, we have
(6.15) $\nabla_{x}u(t,x)=$
$=e^{i\xi^{2}/2(1-\tau)}(1-\tau)^{n/2}\left(i\xi+(1-\tau)\nabla_{\xi}\right)v\left(\tau,\xi\right),$
$(1-\tau)^{n/2}\nabla_{\xi}v(\tau,\xi)$
$=e^{-ix^{2}/(2(1+t))}\left(-ix+(1+t)\nabla_{x}\right)u(t,x).$
Hence,
(6.16) $\int_{\mathbb{R}^{n}}|\nabla_{\xi}v(\tau,\xi)|^{2}d\xi$
$=\int_{\mathbb{R}^{n}}|\left(x+i(1+t)\nabla_{x}\right)u(t,x)|^{2}dx=\|J(1+t)u(t)\|_{L^{2}}^{2}.$
Remark 6.2.
The above relations (6.8) and (6.15) show that
(6.17) $u_{0}\in\mathcal{H}^{1,1}\ \ \mbox{if and only if }\ \ v_{0}(x)=u_{0}(x)e^{-ix^{2}/2}\in\mathcal{H}^{1,1}.$
After the pseudoconformal transformation we find the Cauchy problem (6.6) for $v.$ The precise derivation of energy conservation can be done by following [33], [11].
6.3. Conservation of the pseudoconformal energy
In alternative way we can approximate the initial data with data in $H^{2}$ we can assume that
$v\in C([0,1);H^{2}(\mathbb{R}^{n}))\cap C^{1}([0,1);L^{2}(\mathbb{R}^{n})).$
Multiplying (in $L^{2}$) the equation
(6.18) $(1-\tau)(i\partial_{\tau}+\frac{1}{2}\Delta_{\xi})v=v|v|^{2/n}$
by $\partial_{t}\overline{v}$ and taking the real part, we find
(6.19) $\frac{(1-\tau)}{4}\frac{d}{d\tau}\|\nabla_{\xi}v(\tau)\|_{L^{2}}^{2}+\frac{d}{d\tau}\frac{n}{2(n+1)}\|v(\tau)\|_{L^{2(n+1)/n}}^{2(n+1)/n}=0$
or
(6.20) $\frac{d}{d\tau}\left(\frac{(1-\tau)}{4}\|\nabla v(\tau)\|_{L^{2}}^{2}+\frac{n}{2(n+1)}\|v(\tau)\|_{L^{2(n+1)/n}}^{2(n+1)/n}\right)$
$+\frac{1}{4}\|\nabla u(\tau)\|_{L^{2}}^{2}=0.$
Passing to $t=\tau/(1-\tau),$ we find $\partial_{\tau}=(1+t)^{2}\partial_{t}$, so
(6.21) $\frac{1}{4(1+t)}\frac{d}{dt}\|J(1+t)u(t)\|_{L^{2}}^{2}+\frac{d}{dt}\frac{n}{2(n+1)}\|u(t)\|_{L^{2(n+1)/n}}^{2(n+1)/n}=0$
and
(6.22) $\frac{d}{dt}\left(\frac{1}{4(1+t)}\|J(1+t)u(t)\|_{L^{2}}^{2}+\frac{n}{2(n+1)}\|u(t)\|_{L^{2(n+1)/n}}^{2(n+1)/n}\right)$
$+\frac{1}{4(1+t)^{2}}\|J(1+t)u(t)\|_{L^{2}}^{2}=0.$
Integrating (6.20) from 0 to $\tau$, we find
(6.23) $(1-\tau)\|\nabla v(\tau)\|_{L^{2}}^{2}+\frac{2n}{(n+1)}\|v(\tau)\|_{L^{2(n+1)/n}}^{2(n+1)/n}+\int_{0}^{\tau}\|\nabla v(\theta)\|_{L^{2}}^{2}d\theta=$
$\|\nabla v(0)\|_{L^{2}}^{2}+\frac{2n}{(n+1)}\|v(0)\|_{L^{2(n+1)/n}}^{2(n+1)/n}.$
Integrating (6.22), we obtain
(6.24) $(1+t)^{-1}\|J(1+t)u(t)\|^{2}_{L^{2}}+\frac{2n}{(n+1)}(1+t)\|u(t)\|_{L^{2(n+1)/n}}^{2(n+1)/n}$
$+\int_{0}^{t}\|J(1+\sigma)u(\sigma)\|^{2}_{L^{2}}(1+\sigma)^{-2}d\sigma$
$=\|J(1)u(0)\|^{2}_{L^{2}}+\frac{2n}{(n+1)}\|u(0)\|_{L^{2(n+1)/n}}^{2(n+1)/n}.$
Note that the proof gives the following energy estimate for $v.$
Lemma 6.2.
Let $v$ be a classical solution to (6.6) with initial data in $H^{1}.$ Then for any $\tau\in[0,1)$ we have the identity
(6.25) $(1-\tau)\|\nabla v(\tau)\|^{2}_{L^{2}}+\frac{2n}{n+1}\|v(\tau)\|^{2+2/n}_{L^{2+2/n}}$
$+\int_{0}^{\tau}\|\nabla v(\sigma)\|^{2}_{L^{2}}d\sigma=C_{*},$
$C_{*}=\frac{1}{2}\|\nabla v(0)\|^{2}_{L^{2}}+\frac{2n}{n+1}\|v(0)\|^{2+2/n}_{L^{2+2/n}}.$
Proof of Propostition 6.1.
From (6.24) we conclude that (6.2) is true.
7. Proof of Theorem 2.2
We have the following relation.
(7.1) $\left(u^{+}|\partial_{j}\varphi\right)=\left(\left.u^{+}-U(-t)e^{i\Phi_{t}(u)}u(t)\right|\partial_{j}\varphi\right)$
$-\left(\partial_{j}\left(\left.U(-t)e^{i\Phi_{t}(u)}u(t)\right)\right|\varphi\right).$
Hence, we have the estimates
(7.2) $\left|\left(u^{+}|\partial_{j}\varphi\right)\right|\leq\left\|u^{+}-U(-t)e^{i\Phi_{t}(u)}u(t)\right\|_{L^{2}}\|\partial_{j}\varphi\|_{L^{2}}$
$+\left(\|\partial_{j}\Phi_{t}\|_{L^{2}}\|u(t)\|_{L^{\infty}}+\|\partial_{j}u(t)\|_{L^{2}}\right)\|\varphi\|_{L^{2}}.$
Next, we use the relations
(7.3) $\partial_{j}\left|u\left(\tau,\frac{\tau x}{t}\right)\right|^{2/n}=\frac{2\tau}{nt}\left|u\left(\tau,\frac{\tau x}{t}\right)\right|^{2/n-2}\mathrm{Re}\left(\overline{u\left(\tau,\frac{\tau x}{t}\right)}\ \partial_{j}u\left(\tau,\frac{\tau x}{t}\right)\right)$
$=-\frac{2}{nt}\left|u\left(\tau,\frac{\tau x}{t}\right)\right|^{2/n-2}\mathrm{Im}\left(\overline{u\left(\tau,\frac{\tau x}{t}\right)}\ J_{j}(\tau)u\left(\tau,\frac{\tau x}{t}\right)\right),$
where
(7.4) $J_{j}(\tau)u(\tau,x)=(i\tau\partial_{x_{j}}+x_{j})u(\tau,x).$
Hence, we have
(7.5) $\|\partial_{j}\Phi_{t}\|_{L^{2}}\leq\frac{2}{nt}\int_{1}^{t}\|u(\tau)\|_{L^{\infty}}^{2/n-1}\|J_{j}u\left(\tau,\frac{\tau\ \cdot}{t}\right)\|_{L^{2}}d\tau$
$\leq\frac{2}{nt}\int_{1}^{t}\|u(\tau)\|_{L^{\infty}}^{2/n-1}\left(\frac{\tau}{t}\right)^{-n/2}\|J_{j}(\tau)u\left(\tau\right)\|_{L^{2}}d\tau$
$\leq Ct^{n/2-1}\int_{1}^{t}\tau^{-1}\|J_{j}(\tau)u\left(\tau\right)\|_{L^{2}}d\tau\lesssim t^{(n-1)/2},$
where in the last step, we have used the estimates of Proposition 6.1. From (7.2) we now obtain
(7.6) $\left|\left(u^{+}|\partial_{j}\varphi\right)\right|\leq\left\|u^{+}-U(-t)e^{i\Phi_{t}(u)}u(t)\right\|_{L^{2}}\|\partial_{j}\varphi\|_{L^{2}}$
$+\left(Ct^{-1/2}+2E(u_{0})\right)\|\varphi\|_{L^{2}}$
so taking the limit $t\to+\infty$ we arrive at
(7.7) $\left|\left(u^{+}|\partial_{j}\varphi\right)\right|\lesssim 2E(u_{0})\|\varphi\|_{L^{2}}$
and we can conclude that $u^{+}$ is in $H^{1}.$
In order to show the required estimate (2.5), we approximate $\varphi\in H^{1}$ with a sequence in $H^{2},$ so that
(7.8) $\varphi_{\varepsilon}\in H^{2},\ \ \lim_{\varepsilon\to 0}\|\varphi_{\varepsilon}-\varphi\|_{H^{1}}=0.$
As above, we have the estimates
(7.9) $\left|\left(\nabla\left(U(-t)e^{i\Phi_{t}(u)}u(t)-u^{+}\right)|\nabla\varphi\right)\right|$
$\leq\left(\left\|\nabla U(-t)e^{i\Phi_{t}(u)}u(t)\right\|_{L^{2}}+\|\nabla u^{+}\|_{L^{2}}\right)\|\nabla(\varphi_{\varepsilon}-\varphi)\|_{L^{2}}$
$+\left\|U(-t)e^{i\Phi_{t}(u)}u(t)-u^{+}\right\|_{L^{2}}\|\Delta\varphi_{\varepsilon}\|_{L^{2}}$
$\leq C\left(t^{-1/2}+E(u_{0})\right)\|\nabla(\varphi_{\varepsilon}-\varphi)\|_{L^{2}}$
$+\left\|U(-t)e^{i\Phi_{t}(u)}u(t)-u^{+}\right\|_{L^{2}}\|\Delta\varphi_{\varepsilon}\|_{L^{2}}.$
This leads to the estimate
(7.10) $\limsup_{t\to+\infty}\left|\left(\nabla\left(U(-t)e^{i\Phi_{t}(u)}u(t)-u^{+}\right)|\nabla\varphi\right)\right|\leq CE(u_{0})\|\nabla(\varphi-\varphi_{\varepsilon})\|_{L^{2}}$
so we have (2.5).
8. Weak convergence of modified wave operators
In this section we prove Theorem 2.9 that is a generalization of a result in [21] obtained for the critical Hartree equation in dimensions $n\geq 2.$ Namely, we consider the case $n=2$ and the Cauchy problem
(8.1) $(i\partial_{t}+\frac{1}{2}\Delta_{x})u=u|u|,\ \ t\geq 0,x\in\mathbb{R}^{2},$
$u(0)=u_{0}.$
As in the work [21] we set
(8.2) $\Phi_{t}(u)(x)=\int_{1}^{+\infty}\left|u\left(\tau,\frac{\tau}{t}x\right)\right|d\tau.$
Proof of Theorem 2.9.
As before, (see (4.5) in Section 4) we set
(8.3) $v(t)=M(t)U(-t)\exp\left(i\Phi_{t}(u)\right)\ u(t).$
Lemma 4.1 guarantees that
(8.4) $\mathcal{F}M(t)U(-t)(gu(t))=\tilde{g}\mathcal{F}M(t)U(-t)u(t),\ \ \ \tilde{g}(x)=g(xt).$
Then we have (4.10) and we can refer to Corolary 4.1 that implies
(8.5) $i\frac{d}{dt}v(t)=\mathcal{F}^{-1}\exp\left(i\int_{0}^{t}|u(\tau,\tau x)|d\tau\right)\frac{(-\Delta)}{2t^{2}}\mathcal{F}M(t)U(-t)u(t).$
For any $t>s\geq 1$ and for any $\varphi\in\mathcal{S}(\mathbb{R}^{2})$ we obtain
(8.6) $(v(t)-v(s)|\varphi)=-i\int_{s}^{t}(iv^{\prime}(\tau)|\varphi)d\tau$
$=\frac{i}{2}\int_{s}^{t}\tau^{-2}\left(\mathcal{F}^{-1}\exp\left(i\tilde{\Phi}_{\tau}\right)\Delta\mathcal{F}\mathbb{W}(\tau)|\varphi\right)d\tau$
$=\frac{i}{2}\int_{s}^{t}\tau^{-2}\left(\exp\left(i\tilde{\Phi}_{\tau}\right)\Delta\mathcal{F}\mathbb{W}(\tau)|\mathcal{F}(\varphi)\right)d\tau=$
$=-\frac{i}{2}\int_{s}^{t}\tau^{-2}\left(\exp\left(i\tilde{\Phi}_{\tau}\right)\nabla\mathcal{F}\mathbb{W}(\tau)|\nabla\mathcal{F}(\varphi)\right)d\tau$
$-\frac{i}{2}\int_{s}^{t}\tau^{-2}\left(\nabla\left(\exp\left(i\tilde{\Phi}_{\tau}\right)\right)\nabla\mathcal{F}\mathbb{W}(\tau)|\mathcal{F}(\varphi)\right)d\tau=:I+II,$
where
(8.7) $\mathbb{W}(t)=M(t)U(-t)u(t)$
and
(8.8) $\tilde{\Phi}_{t}(x)=\int_{1}^{t}|u(\tau,\tau x)|d\tau.$
Further we estimate each of the terms $I$ and II.II.
(8.9) $|I|\leq\int_{s}^{t}\tau^{-2}\|J(\tau)u(\tau)\|_{L^{2}}d\tau\|\nabla\mathcal{F}(\varphi)\|_{L^{2}}$
$\leq\left(\int_{s}^{t}\tau^{-2}\right)^{1/2}\left(\int_{s}^{t}\tau^{-2}\|J(\tau)u(\tau)\|^{2}_{L^{2}}d\tau\right)^{1/2}\|\nabla\mathcal{F}(\varphi)\|_{L^{2}}$
$\leq\left(\frac{1}{s}-\frac{1}{t}\right)^{1/2}\left(\int_{s}^{t}\tau^{-2}\|J(\tau)u(\tau)\|^{2}_{L^{2}}d\tau\right)^{1/2}\|\nabla\mathcal{F}(\varphi)\|_{L^{2}}.$
Inequality (5.4) guarantees the smallness of $I.$
The gradient of the phase factor in $II$ is estimated as follows.
(8.10) $\nabla\exp\left(i\tilde{\Phi}_{t}\right)=i\exp\left(i\tilde{\Phi}_{t}\right)\nabla\int_{1}^{t}\left|u\left(\sigma,\sigma x\right)\right|d\sigma$
$=$ $i\exp\left(i\tilde{\Phi}_{t}\right)\mathrm{Re}\int_{1}^{t}\sigma\frac{\overline{u\left(\sigma,\sigma x\right)}}{\left|u\left(\sigma,\sigma x\right)\right|}\nabla u\left(\sigma,\sigma x\right)d\sigma$
$=i\exp\left(i\tilde{\Phi}_{t}\right)\mathrm{Im}\int_{1}^{t}\frac{\overline{u\left(\sigma,\sigma x\right)}}{\left|u\left(\sigma,\sigma x\right)\right|}J(\sigma)u\left(\sigma,\sigma x\right)d\sigma$
and we obtain
(8.11) $\left\|\nabla\exp\left(i\tilde{\Phi}_{\tau}\right)\right\|_{L^{2}}\leq\int_{1}^{\tau}\left\|J(\sigma)u\left(\sigma,\sigma\cdot\right)\right\|_{L^{2}}d\sigma$
$=\int_{1}^{\tau}\sigma^{-1}\left\|J(\sigma)u\left(\sigma\right)\right\|_{L^{2}}d\sigma.$
We have also
(8.12) $\left\|\nabla\mathcal{F}\mathbb{W}(\tau)\right\|_{L^{2}}\lesssim\left\|\nabla\mathcal{F}M(\tau)U(-t)u(\tau)\right\|_{L^{2}}\lesssim\|J(\tau)u(\tau)\|_{L^{2}}.$
In this way, we can obtain
(8.13) $|II|=\left|\frac{1}{2}\int_{s}^{t}\tau^{-2}\left(\nabla\left(\exp\left(i\tilde{\Phi}_{\tau}\right)\right)\nabla\mathcal{F}\mathbb{W}(\tau)|\mathcal{F}(\varphi)\right)d\tau\right|$
$\lesssim\int_{s}^{t}\tau^{-2}\left\|\nabla\left(\exp\left(i\tilde{\Phi}_{\tau}\right)\right)\right\|_{L^{2}}\left\|\nabla\mathcal{F}\mathbb{W}(\tau)\right\|_{L^{2}}\|\|\mathcal{F}(\varphi)\|_{L^{\infty}}d\tau$
$\lesssim\left[\int_{s}^{t}\tau^{-2}\left(\int_{1}^{\tau}\sigma^{-1}\left\|J(\sigma)u\left(\sigma\right)\right\|_{L^{2}}d\sigma\right)\|J(\tau)u(\tau)\|_{L^{2}}d\tau\right]\|\mathcal{F}(\varphi)\|_{L^{\infty}}$
$\lesssim\left(\int_{s}^{t}\tau^{-2}\left(\int_{1}^{\tau}\sigma^{-1}\left\|J(\sigma)u\left(\sigma\right)\right\|_{L^{2}}d\sigma\right)^{2}d\tau\right)^{1/2}$
$\times\left(\int_{s}^{t}\tau^{-2}\|J(\tau)u(\tau)\|^{2}_{L^{2}}d\tau\right)^{1/2}\|\mathcal{F}(\varphi)\|_{L^{\infty}}.$
To this end, we apply the Hardy inequality (5.22) and deduce
(8.14) $|II|\lesssim\left(\int_{s}^{t}\tau^{-2}\|J(\tau)u(\tau)\|^{2}_{L^{2}}d\tau\right)^{1/2}H(t)^{1/2}\|\mathcal{F}(\varphi)\|_{L^{\infty}}$
$=(H(t)-H(s))^{1/2}H(t)^{1/2}\|\mathcal{F}(\varphi)\|_{L^{\infty}},$
where
(8.15) $H(t)=\int_{1}^{t}\tau^{-2}\|J(\tau)u(\tau)\|^{2}_{L^{2}}d\tau.$
Collecting everything, we obtain
(8.16) $\left|(v(t)-v(s)|\varphi)\right|\lesssim\left(\frac{1}{s}-\frac{1}{t}\right)^{1/2}\left(H(t)-H(s)\right)^{1/2}\|\nabla\mathcal{F}(\varphi)\|_{L^{2}}$
$+(H(t)-H(s))^{1/2}H(t)^{1/2}\|\mathcal{F}(\varphi)\|_{L^{\infty}}$
for any $\varphi\in\mathcal{S},$ and for any $t,s$ with $t>s\geq 1.$
It is easy to deduce that for any $\varphi\in L^{2}$ the sequence of linear functionals on $L^{2}$
(8.17) $\Lambda(t)(\varphi)=\left(v(t)|\varphi\right)$
has a pointwise limit as $t\to+\infty.$ In fact, any $\varphi\in L^{2}$ can be approximated by $\varphi_{\varepsilon}\in\mathcal{S}$ in $L^{2}$ so that $\|\varphi-\varphi_{\varepsilon}\|_{L^{2}}\leq\varepsilon.$ So we have
(8.18) $\left|\left(v(t)-v(s)|\varphi\right)\right|$
$\leq\left|\left(v(t)-v(s)|\varphi-\varphi_{\varepsilon}\right)\right|+\left|\left(v(t)-v(s)|\varphi_{\varepsilon}\right)\right|$
$\leq 2\|u_{0}\|_{L^{2}}\varepsilon+\left|\left(v(t)-v(s)|\varphi_{\varepsilon}\right)\right|$
so the estimate (8.16) shows that $\{((v(t)|\varphi),t>1\}$ is a Cauchy sequence and the functionals $\Lambda(t),t>1$ are uniformly bounded on $L^{2}$ due to Banach - Steinhaus theorem so there exists $u_{+}\in L^{2}$ so that
(8.19) $\lim_{t\to+\infty}|((v(t)-u_{+})|\varphi)|=0.$
We can exploit (8.16) to show the convergence in appropriate Banach space. Turning back to the choice $\varphi\in\mathcal{S},$ we see that (8.16) implies also
(8.20) $\left|(v(t)-u_{+}|\varphi)\right|\lesssim\left(\frac{1}{t}\right)^{1/2}\left(H(t)-H_{0}\right)^{1/2}\|\nabla\mathcal{F}(\varphi)\|_{L^{2}}$
$+(H(t)-H_{0})^{1/2}H(t)^{1/2}\|\mathcal{F}(\varphi)\|_{L^{\infty}}.$
where
(8.21) $H_{0}=\lim_{t\to+\infty}H(t)=\int_{1}^{+\infty}\tau^{-2}\|J(\tau)u(\tau)\|^{2}_{L^{2}}d\tau.$
Now we turn to the limit in (2.23).
We take $\varphi\in\mathcal{S}$ and then we have
(8.22) $\left|\left(\left[U(-t)\exp\left(i\Phi_{t}(u)\right)u(t)-u_{+}\right]|\varphi\right)\right|$
$=$ $\left|\left(\left[M(t)U(-t)\exp\left(i\Phi_{t}(u)\right)u(t)-M(t)u_{+}\right]|M(t)\varphi\right)\right|$
$=$ $\left|\left(v(t)-M(t)u_{+}|M(t)\varphi\right)\right|$
$\leq$ $\left|\left(v(t)-M(t)u_{+}|\varphi\right)\right|+\left|\left(v(t)-M(t)u_{+}|M(t)\varphi-\varphi\right)\right|$
$\leq$ $\left|\left(v(t)-u_{+}|\varphi\right)\right|+\left|\left(u_{+}-M(t)u_{+}|\varphi\right)\right|+\left|\left([v(t)-M(t)u_{+}|M(t)\varphi-\varphi\right)\right|$
$\lesssim\left|\left(v(t)-u_{+}|\varphi\right)\right|+(\|u_{0}\|_{L^{2}}+\|u_{+}\|_{L^{2}})\|M(t)\varphi-\varphi\|_{L^{2}}$
$\lesssim\left(\frac{1}{t}\right)^{1/2}\left(H(t)-H_{0}\right)^{1/2}\|\nabla\mathcal{F}(\varphi)\|_{L^{2}}+(H(t)-H_{0})^{1/2}H(t)^{1/2}\|\mathcal{F}(\varphi)\|_{L^{\infty}}$
$+\left(\frac{1}{t}\right)^{1/2}(\|u_{0}\|_{L^{2}}+\|u_{+}\|_{L^{2}})\|\nabla\mathcal{F}(\varphi)\|_{L^{2}}.$
We have continuous embeddings
(8.23) $\dot{B}^{1}_{2,1}\hookrightarrow\dot{B}^{1}_{2,2}=\dot{H}^{1},$
$\dot{B}^{1}_{2,1}\hookrightarrow L^{\infty}$
also the duality relation $(\dot{B}^{1}_{2,1})^{\prime}=\dot{B}^{-1}_{2,{\infty}}$ so we obtain the estimates
(8.24) $\left\|U(-t)\exp\left(i\Phi_{t}(u)\right)u(t)-u_{+}\right\|_{\mathcal{F}\dot{B}^{-1}_{2,\infty}}$
$=$ $\sup_{\varphi\in\mathcal{F}\dot{B}^{1}_{2,1},\|\mathcal{F}\varphi\|_{\dot{B}^{1}_{2,1}}=1}\left|\left(\left[U(-t)\exp\left(i\Phi_{t}(u)\right)u(t)-u_{+}\right]|\varphi\right)\right|$
$\lesssim\left(\frac{1}{\sqrt{t}}+H(t)^{1/2}\right)\left(H(t)-H_{0}\right)^{1/2}+\frac{1}{\sqrt{t}}(\|u_{0}\|_{L^{2}}+\|u_{+}\|_{L^{2}}).$
This completes the proof.
Remark 8.1.
The spaces $\mathcal{F}\dot{B}^{1}_{2,1}$ and $\mathcal{F}\dot{B}^{-1}_{2,\infty}$ correspond to the homogeneous versions of $B_{1}$ and $B_{1}^{*}$ in [1], respectively.
Remark 8.2.
For any Cauchy data in $u_{0}\in\mathcal{H}^{1,1}$ the existence of a weak limit
(8.25) $w-\lim_{t\to{+\infty}}M(t)U(-t)\exp\left(i\Phi_{t}(u)\right)u(t)$
in $L^{2}$ has been proved by Hirata [21].
9. Proof of Theorem 2.10
An important point in the proof of Theorem 2.10 is played by the following.
Proposition 9.1.
For any $z\in\mathbb{C}$ we have
(9.1) $|f(z)|\leq 2|z|^{1+2/n},$
(9.2) $0\leq V(z)\leq 3|z|^{2+2/n},$
(9.3) $W(z)=V(z)+n((1+|z|^{2})^{1/n}-1)-|z|^{2},$
(9.4) $W(z)\leq V(z),$
where $f(z),V(z),W(z)$ are defined in (2.25).
Proof.
For $\lambda>0,$ we put $\varphi(t)=(1+t)^{\lambda}-1,$ $t>0.$ If $0<\lambda<1,$ then
(9.5) $0\leq\varphi(t)=\varphi(t)-\varphi(0)=\int_{0}^{t}\varphi(\tau)d\tau=\lambda\int_{0}^{t}(1+s)^{\lambda-1}ds\leq\lambda t.$
This leads to
(9.6) $|f(z)|=((1+|z|^{2})^{1/n}-1)|z|\leq\frac{|z|^{3}}{n}.$
In particular, if $|z|\leq 1$, then
(9.7) $|f(z)|\leq\frac{1}{n}|z|^{1+2/n}.$
If $|z|\geq 1,$ then we have
(9.8) $|f(z)|\leq(1+|z|^{2})^{1/n}|z|\leq(2|z|^{2})^{1/n}|z|\leq 2|z|^{1+2/n}.$
By (9.7) and (9.8), we have (9.1) .
If $1<\lambda<2,$ then
(9.9) $0\leq\varphi(t)=\varphi^{\prime}(0)t+\int_{0}^{t}(t-s)\varphi^{\prime\prime}(s)ds$
$=\lambda t+\lambda(\lambda-1)\int_{0}^{t}(t-s)(1+s)^{\lambda-2}ds\leq\lambda t+\frac{\lambda(\lambda-1)}{2}t^{2}.$
This implies
(9.10) $0\leq V(z)=\frac{n}{n+1}\left[(1+|z|^{2})^{(n+1)/n}-1)-\frac{n+1}{n}z^{2}\right]$
$\leq\frac{n}{n+1}\frac{n+1}{n}\frac{1}{2n}|z|^{4}=\frac{1}{2n}|z|^{4}.$
In particular, if $|z|\leq 1,$
(9.11) $0\leq V(z)\leq\frac{1}{2n}|z|^{2+2/n}.$
If $|z|\geq 1,$ then we have
(9.12) $0\leq V(z)\leq\frac{n}{n+1}(1+|z|^{2})^{(n+1)/n}$
$\leq\frac{n}{n+1}(2|z|^{2})^{(n+1)/n}\leq\frac{3n}{n+1}|z|^{2+2/n}.$
By (9.10) and (9.12) we have (9.2).
Using the relation
$\langle z\rangle^{2+2/n}-1=|z|^{2}\langle z\rangle^{2/n}+\langle z\rangle^{2/n}-1,$
we find
(9.13) $(n+1)V(z)-n\overline{z}f(z)=(n+1)\left[\frac{n}{n+1}(\langle z\rangle^{2+2/n}-1)-|z|^{2}\right]$
$-n|z|^{2}\left[(\langle z\rangle^{2/n}-1)\right]$
$=\frac{(n+1)n}{n+1}\left(\langle z\rangle^{2/n}-1\right)+\frac{(n+1)n}{n+1}|z|^{2}\langle z\rangle^{2/n}$
$-(n+1)|z|^{2}-n|z|^{2}(\langle z\rangle^{2/n}-1)$
$=n\left(\langle z\rangle^{2/n}-1\right)-|z|^{2}.$
Hence we can write
(9.14) $W(z)=(n+2)V(z)-n\overline{z}f(z)=V(z)+(n+1)V(z)-n\overline{z}f(z)$
$=V(z)+n\left(\langle z\rangle^{2/n}-1\right)-|z|^{2}$
which shows (9.3).
For $\lambda\in(0,1)$ we have
(9.15) $(1+t)^{\lambda}-1-\lambda t=\lambda(\lambda-1)\int_{0}^{t}(t-s)(1+s)^{\lambda-2}ds\leq 0.$
This leads to
(9.16) $n\left(\langle z\rangle^{2/n}-1\right)-|z|^{2}\leq 0$
and (9.4). ∎
Let $u_{0}\in L^{2}(\mathbb{R}^{n}).$ Then there exists a unique global solution $u\in C(\mathbb{R};L^{2})$ in Strichartz class, such that
(9.17) $\|u(t)\|_{L^{2}}=\|u_{0}\|_{L^{2}}$
for any $t\in\mathbb{R}.$
Let $u_{0}\in H^{1}(\mathbb{R}^{n}).$ Then there exists a unique global solution $u\in C(\mathbb{R};H^{1})$ in Strichartz class, such that
(9.18) $E(t)=E(0)$
for any $t\in\mathbb{R},$ where
(9.19) $E(t)=\frac{1}{2}\|\nabla u(t)\|_{L^{2}}^{2}+\int_{\mathbb{R}^{n}}V(u(t,x))dx.$
and $V(z)$ defined in (2.25).
Let $u_{0}\in L^{2}(\mathbb{R}^{n})$ satisfy $xu_{0}\in L^{2}.$ Then the global $L^{2}$- solution $u$ satisfies $xU(-t)u(t)\in L^{2}$ and
(9.20) $\|J(t)u(t)\|_{L^{2}}^{2}+t^{2}\int_{\mathbb{R}^{n}}V(u(t,x))dx=\|xu_{0}\|_{L^{2}}^{2}+\int_{\mathbb{R}^{n}}W(u(t,x))dx$
for any $t\in\mathbb{R}$ and $V,W$ defined in (2.25).
Moreover,
(9.21) $\frac{d}{dt}\left(t^{-1}\left(\|J(t)u(t)\|_{L^{2}}^{2}+t^{2}\int_{\mathbb{R}^{n}}V(u(t,x))dx\right)\right)$
$=-t^{-2}\left(\|J(t)u(t)\|_{L^{2}}^{2}+t^{2}\int_{\mathbb{R}^{n}}V(u(t,x))dx\right)+\int_{\mathbb{R}^{n}}W(u(t,x))dx$
$=-t^{-2}\|J(t)u(t)\|_{L^{2}}^{2}+\int_{\mathbb{R}^{n}}W(u(t,x))dx-\int_{\mathbb{R}^{n}}V(u(t,x))dx$
$\leq-t^{-2}\|J(t)u(t)\|_{L^{2}}^{2}.$
This yields
(9.22) $t^{-1}\|J(t)u(t)\|_{L^{2}}^{2}+t\int_{\mathbb{R}^{n}}V(u(t,x))dx+\int_{1}^{t}s^{-2}\|J(s)u(s)\|_{L^{2}}^{2}ds$
$\leq\|J(1)u(1)\|_{L^{2}}^{2}+\int_{\mathbb{R}^{n}}V(u(1,x))dx$
for any $t\geq 1.$
We define the phase function
(9.23) $\Phi_{t}(u)(x)=\int_{1}^{+\infty}\left[(1+|u(\tau,\tau x/t)|^{2})^{1/n}-1\right]d\tau.$
Proof of Theorem 2.10.
We set
(9.24) $v(t)=M(t)U(-t)\exp\left(i\Phi_{t}(u)\right)\ u(t).$
Lemma 4.1 guarantees that
(9.25) $\mathcal{F}M(t)U(-t)(gu(t))=\tilde{g}\mathcal{F}M(t)U(-t)u(t),\ \ \ \tilde{g}(x)=g(xt).$
Therefore, the Dollard decomposition implies
(9.26) $v(t)=\mathcal{F}^{-1}\mathcal{F}M(t)U(-t)e^{i\Phi_{t}(u)}u(t)=\mathcal{F}^{-1}e^{i\tilde{\Phi}_{t}}\mathcal{F}M(t)U(-t)u(t),$
where
(9.27) $\tilde{\Phi}_{t}(x)=\int_{1}^{t}\left[(1+|u(\tau,\tau x)|^{2})^{1/n}-1\right]d\tau.$
Then we have (4.10) and we can refer to Corolary 4.1 that implies
(9.28) $i\frac{d}{dt}v(t)=\mathcal{F}^{-1}\exp\left(i\tilde{\Phi}_{t}(x)\right)\frac{(-\Delta)}{2t^{2}}\mathcal{F}M(t)U(-t)u(t).$
For any $t>s\geq 1$ and for any $\varphi\in\mathcal{S}(\mathbb{R}^{n})$ we obtain
(9.29) $(v(t)-v(s)|\varphi)=-i\int_{s}^{t}(iv^{\prime}(\tau)|\varphi)d\tau$
$=\frac{i}{2}\int_{s}^{t}\tau^{-2}\left(\mathcal{F}^{-1}\exp\left(i\tilde{\Phi}_{\tau}\right)\Delta\mathcal{F}\mathbb{W}(\tau)|\varphi\right)d\tau$
$=\frac{i}{2}\int_{s}^{t}\tau^{-2}\left(\exp\left(i\tilde{\Phi}_{\tau}\right)\Delta\mathcal{F}\mathbb{W}(\tau)|\mathcal{F}(\varphi)\right)d\tau$
$=-\frac{i}{2}\int_{s}^{t}\tau^{-2}\left(\exp\left(i\tilde{\Phi}_{\tau}\right)\nabla\mathcal{F}\mathbb{W}(\tau)|\nabla\mathcal{F}(\varphi)\right)d\tau$
$-\frac{i}{2}\int_{s}^{t}\tau^{-2}\left(\nabla\left(\exp\left(i\tilde{\Phi}_{\tau}\right)\right)\nabla\mathcal{F}\mathbb{W}(\tau)|\mathcal{F}(\varphi)\right)d\tau=:I+II,$
where
(9.30) $\mathbb{W}(t)=M(t)U(-t)u(t).$
Further we estimate each of the terms $I$ and II.II.
(9.31) $|I|\leq\int_{s}^{t}\tau^{-2}\|J(\tau)u(\tau)\|_{L^{2}}d\tau\|\nabla\mathcal{F}(\varphi)\|_{L^{2}}$
$\leq\left(\int_{s}^{t}\tau^{-2}\right)^{1/2}\left(\int_{s}^{t}\tau^{-2}\|J(\tau)u(\tau)\|^{2}_{L^{2}}d\tau\right)^{1/2}\|\nabla\mathcal{F}(\varphi)\|_{L^{2}}$
$\leq\left(\frac{1}{s}-\frac{1}{t}\right)^{1/2}\left(\int_{s}^{t}\tau^{-2}\|J(\tau)u(\tau)\|^{2}_{L^{2}}d\tau\right)^{1/2}\|\nabla\mathcal{F}(\varphi)\|_{L^{2}}.$
Inequality (5.4) guarantees the smallness of $I.$
The gradient of the phase factor in $II$ is estimated as follows.
(9.32) $\nabla\exp\left(i\tilde{\Phi}_{t}\right)=i\exp\left(i\tilde{\Phi}_{t}\right)\nabla\tilde{\Phi}_{t}$
$=$ $i\frac{2}{n}\exp\left(i\tilde{\Phi}_{t}\right)\int_{1}^{t}\sigma\left(u(\sigma,\sigma x)\right)^{2/n-2}\ \mathrm{Re}\left[\overline{u\left(\sigma,\sigma x\right)}\nabla u\left(\sigma,\sigma x\right)\right]d\sigma$
$=i\frac{2}{n}\exp\left(i\tilde{\Phi}_{t}\right)\int_{1}^{t}\sigma\left(u(\sigma,\sigma x)\right)^{2/n-2}\ \mathrm{Im}\left[\overline{u\left(\sigma,\sigma x\right)}Ju\left(\sigma,\sigma x\right)\right]d\sigma$
and is estimated as
(9.33) $\left\|\nabla\exp\left(i\tilde{\Phi}_{\tau}\right)\right\|_{L^{2}}\leq\frac{2}{n}\int_{1}^{\tau}\left\|J(\sigma)u\left(\sigma,\sigma\cdot\right)\right\|_{L^{2}}d\sigma$
$=\frac{2}{n}\int_{1}^{\tau}\sigma^{-n/2}\left\|J(\sigma)u\left(\sigma\right)\right\|_{L^{2}}d\sigma$
$\lesssim\int_{1}^{\tau}\sigma^{-1}\left\|J(\sigma)u\left(\sigma\right)\right\|_{L^{2}}d\sigma=:f(\tau).$
We have also
(9.34) $\left\|\nabla\mathcal{F}\mathbb{W}(\tau)\right\|_{L^{2}}\lesssim\left\|\nabla\mathcal{F}M(\tau)U(-t)u(\tau)\right\|_{L^{2}}\lesssim\|J(\tau)u(\tau)\|_{L^{2}}$
In this way, we can obtain
(9.35) $|II|=\left|\frac{1}{2}\int_{s}^{t}\tau^{-2}\left(\nabla\left(\exp\left(i\tilde{\Phi}_{\tau}\right)\right)\nabla\mathcal{F}\mathbb{W}(\tau)|\mathcal{F}(\varphi)\right)d\tau\right|$
$\lesssim\int_{s}^{t}\tau^{-2}\left\|\nabla\left(\exp\left(i\tilde{\Phi}_{\tau}\right)\right)\right\|_{L^{2}}\left\|\nabla\mathcal{F}\mathbb{W}(\tau)\right\|_{L^{2}}\|\|\mathcal{F}(\varphi)\|_{L^{\infty}}d\tau$
$\lesssim\left[\int_{s}^{t}\tau^{-2}\left(\int_{1}^{\tau}\sigma^{-1}\left\|J(\sigma)u\left(\sigma\right)\right\|_{L^{2}}d\sigma\right)\|J(\tau)u(\tau)\|_{L^{2}}d\tau\right]\|\mathcal{F}(\varphi)\|_{L^{\infty}}$
$\lesssim\left(\int_{s}^{t}\tau^{-2}\left(\int_{1}^{\tau}\sigma^{-1}\left\|J(\sigma)u\left(\sigma\right)\right\|_{L^{2}}d\sigma\right)^{2}d\tau\right)^{1/2}$
$\times\left(\int_{s}^{t}\tau^{-2}\|J(\tau)u(\tau)\|^{2}_{L^{2}}d\tau\right)^{1/2}\|\mathcal{F}(\varphi)\|_{L^{\infty}}.$
To this end, we apply the Hardy inequality (5.22) and deduce
(9.36) $|II|\lesssim\left(\int_{s}^{t}\tau^{-2}\|J(\tau)u(\tau)\|^{2}_{L^{2}}d\tau\right)^{1/2}H(t)^{1/2}\|\mathcal{F}(\varphi)\|_{L^{\infty}}$
$=(H(t)-H(s))^{1/2}H(t)^{1/2}\|\mathcal{F}(\varphi)\|_{L^{\infty}},$
where
(9.37) $H(t)=\int_{1}^{t}\tau^{-2}\|J(\tau)u(\tau)\|^{2}_{L^{2}}d\tau.$
Collecting everything, we obtain
(9.38) $\left|(v(t)-v(s)|\varphi)\right|\lesssim\left(\frac{1}{s}-\frac{1}{t}\right)^{1/2}\left(H(t)-H(s)\right)^{1/2}\|\nabla\mathcal{F}(\varphi)\|_{L^{2}}$
$+(H(t)-H(s))^{1/2}H(t)^{1/2}\|\mathcal{F}(\varphi)\|_{L^{\infty}}.$
for any $\varphi\in\mathcal{S},$ and for any $t,s$ with $t>s\geq 1.$
It is easy to deduce that for any $\varphi\in L^{2}$ the sequence of linear functionals on $L^{2}$
(9.39) $\Lambda(t)(\varphi)=\left(v(t)|\varphi\right)$
has a pointwise limit as $t\to{+\infty}.$ In fact, any $\varphi\in L^{2}$ can be approximated by $\varphi_{\varepsilon}\in\mathcal{S}$ in $L^{2}$ so that $\|\varphi-\varphi_{\varepsilon}\|_{L^{2}}\leq\varepsilon.$ So we have
(9.40) $\left|\left(v(t)-v(s)|\varphi\right)\right|$
$\leq\left|\left(v(t)-v(s)|\varphi-\varphi_{\varepsilon}\right)\right|+\left|\left(v(t)-v(s)|\varphi_{\varepsilon}\right)\right|$
$\leq 2\|u_{0}\|_{L^{2}}\ \varepsilon+\left|\left(v(t)-v(s)|\varphi_{\varepsilon}\right)\right|$
and the estimate (8.16) shows that $\{((v(t)|\varphi),t>1\}$ is a Cauchy sequence and the functionals $\Lambda(t),t>1$ are uniformly bounded on $L^{2}.$ From Banach - Steinhaus theorem we see that there exists $u_{+}\in L^{2}$ so that
(9.41) $\lim_{t\to{+\infty}}|((v(t)-u_{+}),\varphi)|=0.$
We can exploit (8.16) to show the convergence in appropriate Banach space. Turning back to the choice $\varphi\in\mathcal{S},$ we see that (8.16) implies also
(9.42) $\left|(v(t)-u_{+}|\varphi)\right|\lesssim\left(\frac{1}{t}\right)^{1/2}\left(H(t)-H_{0}\right)^{1/2}\|\nabla\mathcal{F}(\varphi)\|_{L^{2}}$
$+(H(t)-H_{0})^{1/2}H(t)^{1/2}\|\mathcal{F}(\varphi)\|_{L^{\infty}},$
where
(9.43) $H_{0}=\lim_{t\to{+\infty}}H(t)=\int_{1}^{+\infty}\tau^{-2}\|Ju(\tau)\|^{2}_{L^{2}}d\tau.$
Now we turn to the limit in (2.23).
We take $\varphi\in\mathcal{S}$ and then we have
(9.44) $\left|\left(\left[U(-t)\exp\left(i\Phi_{t}(u)\right)u(t)-u_{+}\right]|\varphi\right)\right|$
$=$ $\left|\left(\left[M(t)U(-t)\exp\left(i\Phi_{t}(u)\right)u(t)-M(t)u_{+}\right]|M(t)\varphi\right)\right|$
$=$ $\left|\left(v(t)-M(t)u_{+}|M(t)\varphi\right)\right|$
$\leq$ $\left|\left(v(t)-M(t)u_{+}|\varphi\right)\right|+\left|\left(v(t)-M(t)u_{+}|M(t)\varphi-\varphi\right)\right|$
$\leq$ $\left|\left(v(t)-u_{+}|\varphi\right)\right|+\left|\left(u_{+}-M(t)u_{+}|\varphi\right)\right|+\left|\left([v(t)-M(t)u_{+}|M(t)\varphi-\varphi\right)\right|$
$\lesssim\left|\left(v(t)-u_{+}|\varphi\right)\right|+(\|u_{0}\|_{L^{2}}+\|u_{+}\|_{L^{2}})\|M(t)\varphi-\varphi\|_{L^{2}}$
$\lesssim\left(\frac{1}{t}\right)^{1/2}\left(H(t)-H_{0}\right)^{1/2}\|\nabla\mathcal{F}(\varphi)\|_{L^{2}}+(H(t)-H_{0})^{1/2}H(t)^{1/2}\|\mathcal{F}(\varphi)\|_{L^{\infty}}$
$+\left(\frac{1}{t}\right)^{1/2}(\|u_{0}\|_{L^{2}}+\|u_{+}\|_{L^{2}})\|\nabla\mathcal{F}(\varphi)\|_{L^{2}}.$
We have continuous embeddings
(9.45) $\dot{B}^{n/2}_{2,1}\hookrightarrow\dot{B}^{1}_{2,2}=\dot{H}^{1},$
$\dot{B}^{n/2}_{2,1}\hookrightarrow L^{\infty}$
also the duality relation $(\dot{B}^{n/2}_{2,1})^{\prime}=\dot{B}^{-n/2}_{2,\infty}$ so we obtain the estimates
(9.46) $\left\|U(-t)\exp\left(i\Phi_{t}(u)\right)u(t)-u_{+}\right\|_{\mathcal{F}\dot{B}^{-n/2}_{2,\infty}}$
$=$ $\sup_{\varphi\in\mathcal{F}\dot{B}^{n/2}_{2,1},\|\mathcal{F}\varphi\|_{\dot{B}^{n/2}_{2,1}}=1}\left|\left(\left[U(-t)\exp\left(i\Phi_{t}(u)\right)u(t)-u_{+}\right]|\varphi\right)\right|$
$\lesssim\left(\frac{1}{\sqrt{t}}+H(t)^{1/2}\right)\left(H(t)-H_{0}\right)^{1/2}+\frac{1}{\sqrt{t}}(\|u_{0}\|_{L^{2}}+\|u_{+}\|_{L^{2}}).$
This completes the proof.
10. Virial bound after pseudoconformal transformation
We continue assuming $v$ is a solution to (6.6), replacing $(\tau,\xi)$ by $(t,x)$, we have
(10.1) $(i\partial_{t}+\frac{1}{2}\Delta_{x})v=\frac{v(t)|v(t)|^{2/n}}{1-t},\ \ \ t\in[0,1);$
$v(0)=v_{0}\in\mathcal{H}^{1,1}.$
Lemma 10.1.
For any $n\geq 1$ and any $R>0$ there exists a constant $C=C(R)>0$ so that if $v$ solves the problem (10.1) with initial data satisfying
(10.2) $\|v_{0}\|_{H^{1}}+\||x|v_{0}\|_{L^{2}}\leq R,$
then $v$ satisfies the inequality
(10.3) $\int_{\mathbb{R}^{n}}|x|^{2}|v(t,x)|^{2}dx\leq C(R).$
Proof.
It is sufficient to prove (10.3) assuming $v_{0}\in\mathcal{H}^{N,1}$ with $N>n/2+2.$ Then $v\in C([0,1);\mathcal{H}^{N,1}).$ We have the following continuity equation, obtained by multiplying (10.1) by $\overline{v}$ and taking the imaginary part
(10.4) $\partial_{t}|v(t,x)|^{2}+\nabla_{x}\left[\mathrm{Im}\left(\nabla_{x}v(t,x)\overline{v(t,x)}\right)\right]=0.$
Now we introduce the following momentum associated with $v$
(10.5) $M_{v}(t)=\int_{\mathbb{R}^{n}}|x|^{2}|v(t,x)|^{2}dx.$
Multiplying the equation (10.4) by the weight $|x|^{2}$ and integrating by parts in $x$ we find
(10.6) $M_{v}(t)\leq M_{v}(0)+2\int_{0}^{t}\sqrt{M_{v}(\tau)}\|\nabla_{x}v(\tau)\|_{L^{2}}d\tau.$
From this estimate we can get
(10.7) $M_{v}(t)\leq M_{v}(0)$
$+2\left(\sup_{0\leq\tau\leq t}M_{v}(\tau)\right)^{1/2}\sup_{0\leq\tau\leq t}(1-\tau)^{1/2}\|\nabla_{x}v(\tau)\|_{L^{2}}\int_{0}^{t}(1-\tau)^{-1/2}d\tau$
$\leq M_{v}(0)+\frac{1}{2}\left(\sup_{0\leq\tau\leq t}M_{v}(\tau)\right)+2\left(\sup_{0\leq\tau\leq t}(1-\tau)^{1/2}\|\nabla_{x}v(\tau)\|_{L^{2}}\right)^{2}.$
Recall the energy identity (6.25) that implies
(10.8) $\sup_{0\leq\tau\leq t}(1-\tau)^{1/2}\|\nabla_{x}v(\tau)\|_{L^{2}}\leq C_{1}(R)$
and we deduce
(10.9) $\sup_{0\leq\tau\leq t}M_{v}(\tau)\leq M_{v}(0)+\frac{1}{2}\left(\sup_{0\leq\tau\leq t}M_{v}(\tau)\right)+2C_{1}(R)^{2}$
so
(10.10) $M_{v}(t)\leq 2M_{v}(0)+4C_{1}(R)^{2}.$
This completes the proof.
11. Transformation of NLS to nonlinear problem for inverted harmonic oscillator
In this section, we continue to study the Cauchy problem (10.1), so that we are looking for a transformation of the NLS in (10.1) to a new nonlinear evolution problem associated with appropriate Hamiltonian such that the uniform bounds
(11.1) $\|v(t)\|_{L^{2}}=const,$
$(1-t)^{1/2}\|\nabla_{x}v(t)\|_{L^{2}}\lesssim 1$
can be interpreted as mass and kinetic energy for the new Hamiltonian. We make the following substitution $(t,x,v)\mapsto(s,y,Y)$ defined by
(11.2) $s=t,\ \ y=\frac{x}{(1-t)^{1/2}},$
$Y(s,y)=e^{iy^{2}/4}(1-s)^{n/4}v(s,y(1-s)^{1/2})$
with inverse substitution given by
(11.3) $t=s,\ x=y(1-s)^{1/2},$
$v(t,x)=e^{-\frac{ix^{2}}{4(1-t)}}(1-t)^{-n/4}Y\left(t,\frac{x}{\sqrt{1-t}}\right).$
Lemma 11.1.
The transformation (11.2), namely
(11.4) $Y(s,y)=e^{iy^{2}/4}(1-s)^{n/4}v(s,y(1-s)^{1/2})$
transforms the equation
(11.5) $\left(i\partial_{t}+\frac{\Delta_{x}}{2}\right)v(t,x)=\frac{v(t,x)|v(t,x)|^{2/n}}{1-t},$
into
(11.6) $i(1-s)\partial_{s}Y(s,y)-\mathcal{H}_{1/2}Y(s,y)=\frac{Y(s,y)|Y(s,y)|^{2/n}}{(1-s)^{1/2}},$
where $\mathcal{H}_{\omega}$ is inverted harmonic oscillator Hamiltonian
(11.7) $\mathcal{H}_{\omega}=-\frac{1}{2}\Delta_{y}-\frac{\omega^{2}}{2}y^{2}.$
Proof.
A direct substitution
(11.8) $v(t,x)=e^{-\frac{ix^{2}}{4(1-t)}}(1-t)^{-n/4}\ Y\left(t,\frac{x}{\sqrt{1-t}}\right)$
gives
(11.9) $(1-t)^{1+n/4}e^{\frac{ix^{2}}{4(1-t)}}\left(i\partial_{t}+\frac{\Delta_{x}}{2}\right)v(t,x)$
$=\left[\left(i(1-t)\partial_{t}Y+\frac{\Delta_{y}}{2}Y\right)\left(t,\frac{x}{\sqrt{1-t}}\right)+\frac{x^{2}}{8(1-t)}Y\left(t,\frac{x}{\sqrt{1-t}}\right)\right].$
On the other hand, we have
(11.10) $e^{\frac{ix^{2}}{4(1-t)}}(1-t)^{1+n/4}\frac{v(t,x)|v(t,x)|^{2/n}}{1-t}$
$=(1-t)^{-1/2}Y\left(t,\frac{x}{\sqrt{1-t}}\right)\left|Y\left(t,\frac{x}{\sqrt{1-t}}\right)\right|^{2/n}.$
Hence,
(11.11) $e^{\frac{ix^{2}}{4(1-t)}}(1-t)^{1+n/4}\left[\left(i\partial_{t}+\frac{\Delta_{x}}{2}\right)v(t,x)-\frac{v(t,x)|v(t,x)|^{2/n}}{1-t}\right]$
$=\left[\left(i(1-t)\partial_{t}Y+\frac{\Delta_{y}}{2}Y\right)\left(t,\frac{x}{\sqrt{1-t}}\right)+\frac{x^{2}}{8(1-t)}Y\left(t,\frac{x}{\sqrt{1-t}}\right)\right]$
$-(1-t)^{-1/2}\ Y\left(t,\frac{x}{\sqrt{1-t}}\right)\left|Y\left(t,\frac{x}{\sqrt{1-t}}\right)\right|^{2/n}.$
Lemma 11.2.
Let $v(t,x)$ be a solution to the Cauchy problem (10.1) and $Y(s,y)$ is defined by (11.2). Then we have the following properties.
  1. i)
    We have the relations
    (11.12) $\|Y(t)\|^{2}_{L^{2}}=\|v(t)\|_{L^{2}}^{2},$
    $\|Y(t)\|^{2(n+1)/n}_{L^{2(n+1)/n}}=(1-t)^{1/2}\|v(t)\|^{2(n+1)/n}_{L^{2(n+1)/n}}.$
  2. ii)
    We have the uniform bounds
    (11.13) $\left\|\left(\nabla_{y}-\frac{iy}{2}\right)Y(t)\right\|^{2}_{L^{2}}\lesssim 1,$
    $(1-t)\||y|Y(t)\|^{2}_{L^{2}}=\||x|v(t)\|^{2}_{L^{2}}\lesssim 1.$
Proof.
We use (11.4) and can write
(11.14) $\|Y(t)\|^{2}_{L^{2}}=\int_{\mathbb{R}^{n}}|v(t,y(1-t)^{1/2})|^{2}(1-t)^{n/2}dy=\|v(t)\|^{2}_{L^{2}},$
and
(11.15) $\|Y(t)\|^{q}_{L^{q}}=\int_{\mathbb{R}^{n}}|v(t,y(1-t)^{1/2})|^{2}(1-t)^{qn/4}dy$
$=(1-t)^{qn/4-n/2}\|v(t)\|^{q}_{L^{q}},$
with $q=2(n+1)/n$ we find
(11.16) $\|Y(t)\|^{2(n+1)/n}_{L^{2(n+1)/n}}=(1-t)^{2(n+1)/4-n/2}\|v(t)\|^{2(n+1)/n}_{L^{2(n+1)/n}}$
$=(1-t)^{1/2}\|v(t)\|^{2(n+1)/n}_{L^{2(n+1)/n}}.$
This proves (11.12).
Now we use the definition (11.4) of $Y$ and find
(11.17) $\left(\nabla_{y}-\frac{iy}{2}\right)Y(s)$
$=\left(\nabla_{y}-\frac{iy}{2}\right)\left(e^{iy^{2}/4}(1-s)^{n/4}v(s,y(1-s)^{1/2})\right)$
$=e^{iy^{2}/4}(1-s)^{n/4+1/2}\nabla v(s,y(1-s)^{1/2}).$
Hence,
(11.18) $\left\|\left(\nabla_{y}-\frac{iy}{2}\right)Y(s)\right\|^{2}_{L^{2}}$
$=\int_{\mathbb{R}^{n}}(1-s)^{n/2+1}|\nabla v(s,y(1-s)^{1/2})|^{2}dy$
$=(1-s)\|\nabla v(s)\|_{L^{2}}^{2}.$
Now we apply the energy estimate
(11.19) $(1-s)\|\nabla v(s)\|_{L^{2}}^{2}\lesssim 1$
following from the energy identity (6.25) in Lemma 6.2 and deduce the corresponding inequality in (11.13).
Finally, we have
(11.20) $(1-t)\||y|Y(t)\|^{2}_{L^{2}}$
$=\int_{\mathbb{R}^{n}}(1-t)|y|^{2}|v(t,y(1-t)^{1/2})|^{2}(1-t)^{n/2}dy$
$=\||x|v(t)\|^{2}_{L^{2}}.$
Now we can use Lemma 10.1 and see that
(11.21) $(1-t)\||y|Y(t)\|^{2}_{L^{2}}\lesssim 1.$
This completes the proof of (11.13) and the Lemma. ∎
12. Virial estimates for inverted harmonic oscillator
12.1. On linear inverted harmonic oscillator
We have the following information for the Hamiltonian known also a inverted harmonic oscillator
(12.1) $\mathcal{H}_{\omega}=-\frac{1}{2}\Delta_{y}-\frac{\omega^{2}}{2}y^{2},\ \omega>0:$
  1. i)
    The operator $\mathcal{H}_{\omega}$ with domain
    (12.2) $\Sigma_{2}=\{f\in H^{2},|y|^{2}f\in L^{2}\}$
    is self - adjoint with absolutely continuous spectrum
    (12.3) $\sigma(\mathcal{H}_{\omega})=\sigma_{ac}(\mathcal{H}_{\omega})=\mathbb{R}.$
  2. ii)
    The operator has no resonances. Namely, the result in [23] guarantees that any solution to
    (12.4) $(\mathcal{H}_{\omega}-\lambda)U=0$
    with
    $U\in\langle x\rangle^{s}L^{2},\ \ s<\frac{1}{2}$
    has to be the trivial one $U=0.$
  3. iii)
    We have limiting absorption principle and $L^{2}$ scattering for appropriate perturbations of $\mathcal{H}_{\omega}$ (see [3], [23]).
12.2. Virial bound for inverted harmonic oscillator
We assume that $Y(s,y),s\in[0,1),y\in\mathbb{R}^{n}$ is a solution to the Cauchy problem
(12.5) $i(1-s)\partial_{s}Y(s,y)-\mathcal{H}_{1/2}Y(s,y)=\frac{Y(s,y)|Y(s,y)|^{2/n}}{(1-s)^{1/2}},$
$Y(0)=Y_{0}\in\mathcal{H}^{1,1}.$
Lemma 12.1.
(Morawetz identity: first step) We assume $Y$ is a solution to (12.5). Then we have
(12.6) $\frac{(1-t)}{2}M_{Y}^{\prime}(t)=\frac{1}{2}\mathrm{Im}\left(\left.LY(t)\right|\overline{Y(t)}\right),$
where
(12.7) $\left(\left.f\right|g\right)=\int_{\mathbb{R}^{n}}f(y)\overline{g(y)}dy,$
(12.8) $M_{Y}(t)=\int_{\mathbb{R}^{n}}y^{2}|Y(t,y)|^{2}dy$
and
(12.9) $L=y\cdot\nabla_{y}+\nabla_{y}\cdot y$
is the dilation generator.
Proof.
Again as in the proof of Lemma 10.1 it is sufficient to prove (12.6) assuming $Y_{0}\in\mathcal{H}^{N,1}$ with $N>n/2+2.$ Then $Y\in C([0,1);\mathcal{H}^{N,1}).$
We have the following continuity equation obtained by multiplying (11.6) by $\overline{Y}$ and taking the imaginary part
(12.10) $\frac{(1-t)}{2}\partial_{t}|Y(t,y)|^{2}+\frac{1}{2}\nabla_{y}\left[\mathrm{Im}\left(\nabla_{y}Y(t,y)\overline{Y(t,y)}\right)\right]=0.$
Multiplying by $y^{2}$ and integrating in $y$, we find
(12.11) $0=\frac{(1-t)}{2}\partial_{t}\left(\int_{\mathbb{R}^{n}}y^{2}|Y(t,y)|^{2}dy\right)-\left(\int_{\mathbb{R}^{n}}\mathrm{Im}\left(y\cdot\nabla_{y}Y(t,y)\overline{Y(t,y)}\right)dy\right)$
$=\frac{(1-t)}{2}\partial_{t}\left(\int_{\mathbb{R}^{n}}y^{2}|Y(t,y)|^{2}dy\right)-\frac{1}{2}\int_{\mathbb{R}^{n}}\mathrm{Im}\left(y\cdot\nabla_{y}Y(t,y)\overline{Y(t,y)}\right)dy$
$+\frac{1}{2}\int_{\mathbb{R}^{n}}\mathrm{Im}\left(Y(t,y)\overline{y\cdot\nabla_{y}Y(t,y)}\right)dy$
$=\frac{(1-t)}{2}\partial_{t}\left(\int_{\mathbb{R}^{n}}y^{2}|Y(t,y)|^{2}dy\right)-\frac{1}{2}\mathrm{Im}\left(\left.\left(y\cdot\nabla_{y}+\nabla_{y}\cdot y\right)Y(t)\right|Y(t)\right).$
Lemma 12.2.
There exists a constant $C_{0}\in\mathbb{R}$ so that we have the asymptotic expansion
(12.12) $M_{Y}(t)=\frac{C_{0}}{1-t}+rem(t),$
$|rem(t)|\leq C(1-t)^{-1/2},$
$|rem^{\prime}(t)|\leq C(1-t)^{-3/2}.$
Proof.
We set
(12.13) $Z(t)=(1-t)M_{Y}(t)$
and applying Lemma 12.1, we have
(12.14) $\frac{1}{2}Z^{\prime}(t)=\frac{1-t}{2}M_{Y}^{\prime}(t)-\frac{1}{2}M_{Y}(t),$
$=\frac{1}{2}\mathrm{Im}\left(\int_{\mathbb{R}^{n}}\left[(y\cdot\nabla_{y}+\nabla_{y}\cdot y)Y(t,y)\right]\,\overline{Y(t,y)}dy\right)$
$-\frac{1}{2}\mathrm{Im}\left(\int_{\mathbb{R}^{n}}\left[(iy\cdot y+iy\cdot y)Y(t,y)\right]\ \overline{Y(t,y)}dy\right)$
$=\frac{1}{2}\mathrm{Im}\left(\int_{\mathbb{R}^{n}}\left[(y\cdot(\nabla_{y}-iy)+(\nabla_{y}-iy)\cdot y)Y(t,y)\right]\ \overline{Y(t,y)}dy\right).$
Therefore, we use the estimates of Lemma 11.2 and find
(12.15) $|Z^{\prime}(t)|\leq C\sqrt{M_{Y}(t)}\leq C(1-t)^{-1/2}.$
We know that $Z(t),t\in[0,1)$ is bounded. Therefore, there exists $t_{k}\nearrow 1$ so that
(12.16) $\lim_{k\to\infty}Z(t_{k})=C_{0}.$
From this and the estimate of $Z^{\prime}$ it is easy to see that $\lim_{t\nearrow 1}Z(t)$ exists, we denote this limit by $C_{0}.$ Then we can write
(12.17) $Z(t)=C_{0}-\int_{t}^{1}Z^{\prime}(\tau)d\tau,$
and then from
(12.18) $rem(t)=-\frac{1}{1-t}\int_{t}^{1}Z^{\prime}(\tau)d\tau,$
we get
(12.19) $|rem(t)|\leq C(1-t)^{-1/2}.$
From (12.18) we can estimate the derivative of $rem(t)$ by using (12.15), so we can write
(12.20) $|rem^{\prime}(t)|\leq\left|\frac{1}{(1-t)^{2}}\int_{t}^{1}Z^{\prime}(\tau)d\tau\right|$
$+\left|\frac{1}{1-t}Z^{\prime}(t)d\tau\right|\lesssim(1-t)^{-3/2}.$
Lemma 12.3.
(Morawetz identity: second step) If $Y$ is a solution to
(12.21) $i(1-t)\partial_{t}Y(t,y)-\mathcal{H}_{\omega}Y(t,y)=\frac{Y(t,y)|Y(t,y)|^{2/n}}{(1-t)^{1/2}},$
where $\mathcal{H}_{\omega},\omega>0$ is inverted harmonic oscillator Hamiltonian
(12.22) $\mathcal{H}_{\omega}=-\frac{1}{2}\Delta_{y}-\frac{\omega^{2}}{2}y^{2},$
then we have
(12.23) $\frac{1}{4}\left((1-t)\frac{d}{dt}\right)^{2}M_{Y}(t)=\|\nabla_{y}Y(t)\|_{L^{2}}^{2}+\omega^{2}M_{Y}(t)$
$+\frac{n}{(n+1)\sqrt{1-t}}\|Y(t)\|_{L^{2+2/n}}^{2+2/n}.$
Proof.
We use the dilation generator $L$ defined in (12.9) which is commonly used in virial and Morawetz identities. Then, we have the commutator relation
(12.24) $\left[\mathcal{H}_{\omega},L\right]=-\left[\frac{1}{4}\Delta+\frac{1}{4}\omega^{2}|y|^{2},y\cdot\nabla_{y}+\nabla_{y}\cdot y\right]=-\Delta+\omega^{2}y^{2}.$
Here we use the commutator relations
(12.25) $\left[\Delta,(y\cdot\nabla_{y}+\nabla_{y}\cdot y)\right]=4\Delta,$
$\left[y^{2},(y\cdot\nabla_{y}+\nabla_{y}\cdot y)\right]=-4y^{2}.$
Using equation (11.6), we apply the operator $L=y\cdot\nabla_{y}+\nabla_{y}\cdot y$ and find
(12.26) $i(1-t)\partial_{t}LY(t,y)-L\mathcal{H}_{\omega}Y(t,y)=L\left(\frac{Y(t,y)|Y(t,y)|^{2/n}}{(1-t)^{1/2}}\right)$
We take the real part of the inner product in $L^{2}$ of the equation in (12.26) and $Y$ and find
(12.27) $-\frac{1-t}{2}\frac{d}{dt}\mathrm{Im}\left(\left.LY(t)\right|Y(t)\right)-\mathrm{Re}\left(\left.LH_{\omega}Y(t)\right|Y(t)\right)$
$=\mathrm{Re}\left(\left.L\left(\frac{Y(t)|Y(t)|^{2/n}}{\sqrt{1-t}}\right)\right|Y(t)\right),$
since
(12.28) $\frac{d}{dt}\mathrm{Im}\left(\left.LY(t)\right|Y(t)\right)=\mathrm{Im}\left(\left.L\partial_{t}Y(t)\right|Y(t)\right)$
$+\mathrm{Im}\left(\left.LY(t)\right|\partial_{t}Y(t)\right)=2\mathrm{Im}\left(\left.L\partial_{t}Y(t)\right|Y(t)\right).$
The second term in (12.27) can be transformed by using (12.24) as follows
(12.29) $-\mathrm{Re}\left(\left.L\mathcal{H}_{\omega}Y(t)\right|Y(t)\right)$
$=-\frac{1}{2}\left(\left.L\mathcal{H}_{\omega}Y(t)\right|Y(t)\right)-\frac{1}{2}\overline{\left(\left.L\mathcal{H}_{\omega}Y(t)\right|Y(t)\right)}$
$=-\frac{1}{2}\left(\left.(L\mathcal{H}_{\omega}-\mathcal{H}_{\omega}L)Y(t)\right|Y(t)\right)$
$=-\frac{1}{2}\left(\left.[L,\mathcal{H}_{\omega}]Y(t)\right|Y(t)\right)$
$=-\frac{1}{4}\left(\left.[(\Delta+\omega^{2}y^{2}),(y\cdot\nabla_{y}+\nabla_{y}\cdot y)]Y(t))\right|Y(t)\right)$
$=\left(\left.(-\Delta+\omega^{2}y^{2})Y(t)\right|Y(t)\right).$
The last term in (12.27) can be transformed by using the following relation
(12.30) $\mathrm{Re}\left(\left(Y(t,y)|Y(t,y)|^{2/n}\right)\nabla_{y}\overline{Y(t,y)}\right)=\frac{n}{2(n+1)}\nabla_{y}|Y(t,y)|^{2+2/n}$
so we have
(12.31) $\mathrm{Re}\left(\left.L\left(Y(t,y)|Y(t,y)|^{2/n}\right)\right|Y(t)\right)$
$=2\mathrm{Re}\left(\int_{\mathbb{R}^{n}}\left[\nabla_{y}\cdot y\left(Y(t,y)|Y(t,y)|^{2/n}\right)\right]\overline{Y(t,y)}dy\right)-n\|Y(t)\|_{L^{2+2/n}}^{2+2/n}$
$=-2\left(\int_{\mathbb{R}^{n}}y\cdot\mathrm{Re}\left(Y(t,y)|Y(t,y)|^{2/n}\right)\nabla_{y}\overline{Y(t,y)}dy\right)-n\|Y(t)\|_{L^{2+2/n}}^{2+2/n}$
$=-\frac{n}{n+1}\int_{\mathbb{R}^{n}}y\cdot\nabla_{y}(|Y(t,y)|^{2+2/n})dy-n\|Y(t)\|_{L^{2+2/n}}^{2+2/n}$
$=\left(\frac{n^{2}}{n+1}-n\right)\|Y(t)\|_{L^{2+2/n}}^{2+2/n}=-\frac{n}{n+1}\|Y(t)\|_{L^{2+2/n}}^{2+2/n}.$
Now (12.27) leads to
(12.32) $\frac{1-t}{2}\frac{d}{dt}\mathrm{Im}\left(\int_{\mathbb{R}^{n}}LY(t,y)\overline{Y(t,y)}dy\right)$
$=\int_{\mathbb{R}^{n}}((-\Delta+\omega^{2}y^{2})Y(t,y))\overline{Y(t,y)}dy+\frac{n}{(n+1)\sqrt{1-t}}\|Y(t)\|_{L^{2+2/n}}^{2+2/n}$
$=\|\nabla_{y}Y(t)\|_{L^{2}}^{2}+\omega^{2}M_{Y}(t)+\frac{n}{(n+1)\sqrt{1-t}}\|Y(t)\|_{L^{2+2/n}}^{2+2/n}.$
Turning back to (12.6), we obtain (12.23).
Lemma 12.4.
The constant $C_{0}$ from Lemma 12.2 is identically zero.
Proof.
We estimate the right side of (12.23), by using the estimates of Lemma 11.2.
(12.33) $\|\nabla_{y}Y(t)\|_{L^{2}}^{2}+\omega^{2}M_{Y}(t)+\frac{n}{(n+1)\sqrt{1-t}}\|Y(t)\|_{L^{2+2/n}}^{2+2/n}$
$=2\omega^{2}M_{Y}(t)+O(\sqrt{M_{Y}(t)})+O(1).$
We further use the quantity $Z(t)$ introduced in (12.13). We use the simple relation
(12.34) $(1-t)\frac{d}{dt}M_{Y}(t)=Z^{\prime}(t)+M_{Y}(t)$
and find for the left side of (12.23)
(12.35) $\frac{1}{4}\left((1-t)\frac{d}{dt}\right)^{2}M_{Y}(t)=\frac{1}{4}\left((1-t)\frac{d}{dt}\right)(Z^{\prime}(t)+M_{Y}(t))$
$=\frac{1}{4}(1-t)Z^{\prime\prime}(t)+\frac{1}{4}Z^{\prime}(t)+\frac{1}{4}M_{Y}(t).$
From (12.23) we find with $\omega=1/2$
(12.36) $\frac{1}{4}(1-t)Z^{\prime\prime}(t)=\frac{1}{4}M_{Y}(t)+O((1-t)^{-1/2}),$
where we have used estimate (12.15). Now we are in position to use the asymptotic expansion (12.12) for $M_{Y}(t)$ and deduce
(12.37) $Z^{\prime\prime}(t)=\frac{C_{0}}{(1-t)^{2}}+O((1-t)^{-3/2}).$
Integrating, we get
(12.38) $Z^{\prime}(t)=Z^{\prime}(0)+\int_{0}^{t}\left(\frac{C_{0}}{(1-\tau)^{2}}+O((1-\tau)^{-3/2})\right)d\tau$
$=\frac{C_{0}}{1-t}+O((1-t)^{-1/2}).$
This representation and (12.15) imply $C_{0}=0.$
Lemma 12.5.
We have
(12.39) $M_{Y}(t)\lesssim 1,\ \forall t\in[0,1).$
Proof.
We use the inequality (12.15) so we can write
(12.40) $|Z^{\prime}(t)|\leq C\sqrt{Z(t)}(1-t)^{-1/2}.$
Integrating with $0<t_{1}<t_{2}<1$, we get
(12.41) $\sqrt{Z(t_{2})}-\sqrt{Z(t_{1})}\geq-C\left(\sqrt{1-t_{1}}-\sqrt{1-t_{2}}\right).$
We know that $Z(t_{2})\to 0$ when $t_{2}\to 1,$ so letting $t_{2}\to 0$ we get
(12.42) $C\sqrt{1-t_{1}}\geq\sqrt{Z(t_{1})}$
and we arrive at
(12.43) $Z(t_{1})\leq C(1-t_{1}).$
This completes the proof.
From the boundedness of $M_{y}(t)$ and relation (11.13) in Lemma 11.2 we obtain the estimate
Lemma 12.6.
If $v$ solves the Cauchy problem (10.1), i.e.
(12.44) $\left(i\partial_{t}+\frac{\Delta_{x}}{2}\right)v(t,x)=\frac{v(t,x)|v(t,x)|^{2/n}}{1-t},\ \ t\in(0,1),x\in\mathbb{R}^{n}$
$v(0,x)=v_{0}(x)\in\mathcal{H}^{1,1},$
then we have the uniform bound
(12.45) $\int_{\mathbb{R}^{n}}|x|^{2}|v(t,x)|^{2}dx\leq C(1-t).$
12.3. Proofs of Theorems 2.6, 2.7 and 2.8.
Proof of Theorem 2.6 .
The proof for $n=1$ is based on the following simple inequality
(12.46) $|f(x)|^{2}\leq\frac{2}{x}\left(\int_{x}^{\infty}|f^{\prime}(y)|^{2}dy\right)^{1/2}\left(\int_{x}^{\infty}|f(y)|^{2}y^{2}dy\right)^{1/2},$
where $x>0.$
The key point is the following uniform bound (see Lemma 12.6)
(12.47) $\int_{\mathbb{R}^{n}}|x|^{2}|v(t,x)|^{2}dx\leq C(1-t).$
The energy estimate (see Lemma 6.2)
(12.48) $\int_{\mathbb{R}^{n}}|\nabla_{x}v(t,x)|^{2}dx\leq C(1-t)^{-1}.$
combined with (12.46) and (12.47) imply (2.15) for $x>1.$ Similarly we proceed for x<1.x<-1. Finally, for $|x|\leq 1$ we use the translation invariance (in $x$) of the equation
$(i\partial_{t}+\frac{1}{2}\Delta_{x})v=\frac{v|v|^{2/n}}{1-t}$
and complete the proof of Theorem 2.6.
Proof of Theorem 2.7 .
For Theorem 2.7 we use Caffarelli - Kohn - Nirenberg inequality ( see (13.1) that is obtained as partial case of more general inequality obtained in [4]).
(12.49) $\left\||x|^{1/2}u\right\|^{2}_{L^{q}}\lesssim\||\nabla u\|_{L^{2}}\||x|u\|_{L^{2}}.$
Using the translation invariance of the Schrödinger equation for $v$, we can obtain the following variant of (12.47), namely
(12.50) $\int_{\mathbb{R}^{n}}|x-x_{0}|^{2}|v(t,x)|^{2}dx\lesssim(1-t)$
provided $|x_{0}|=1.$ In this way we get
(12.51) $\left\||x-x_{0}|^{1/2}u\right\|^{2}_{L^{q}}\lesssim\||\nabla u\|_{L^{2}}\||x-x_{0}||u\|_{L^{2}}\lesssim 1.$
From
(12.52) $|x-x_{0}|^{1/2}+|x|^{1/2}\gtrsim 1$
we complete the proof. ∎
Proof of Theorem 2.8 .
For Theorem 2.8 we use the following variant of (12.46).
(12.53) $r^{n-1}|f(r)|^{2}$ $\leq\frac{2}{r}\left(\int_{r}^{\infty}|f^{\prime}(\tau)|^{2}\tau^{n-1}d\tau\right)^{1/2}\left(\int_{r}^{\infty}|f(\tau)|^{2}\tau^{2}\tau^{n-1}d\tau\right)^{1/2}$
$+\frac{n-1}{r}\int_{r}^{\infty}|f(\tau)|^{2}\tau^{n-1}d\tau.$
13. Caffarelli - Kohn - Nirenberg inequality
We recall the classical Caffarelli - Kohn - Nirenberg inequality (see [4]) involving parameters
$p_{1},p_{2}\geqslant 1,\quad q>0,\quad 0\leqslant a\leqslant 1$
$\frac{1}{p_{1}}+\frac{\alpha}{n},\quad\frac{1}{p_{2}}+\frac{\beta}{n},\quad\frac{1}{q}+\frac{\gamma}{n}>0,$
where
$\gamma=a\sigma+(1-a)\beta.$
If we require $a\in(0,1)$, then the main result in [4] implies the following.
Theorem 13.1.
If
$0\leqslant\alpha-\sigma$
and
$\frac{1}{q}+\frac{\gamma}{n}=a\left(\frac{1}{p_{1}}+\frac{\alpha-1}{n}\right)+(1-a)\left(\frac{1}{p_{2}}+\frac{\beta}{n}\right),$
then there exists a positive constant $C$ such that the following inequality holds for all $u\in C_{0}^{\infty}\left(\mathbb{R}^{n}\right)$
$\left\||x|^{\gamma}u\right\|_{L^{q}}\leqslant C\||x|^{\alpha}\nabla u\|_{L^{p_{1}}}^{a}\||x|^{\beta}u\|_{L^{p_{2}}}^{1-a}$
In the special case $a=1/2,$ $p_{1}=p_{2}=2$, $\alpha=0,$ $\beta=1$ we take $\sigma=\alpha=0$ and deduce $\gamma=1/2$ and
(13.1) $\left\||x|^{1/2}u\right\|^{2}_{L^{q}}\lesssim\|\nabla u\|_{L^{2}}\||x|u\|_{L^{2}}$
provided
(13.2) $\frac{1}{q}+\frac{1}{2n}=\frac{1}{2}\left(\frac{1}{2}-\frac{1}{n}\right)+\frac{1}{2}\left(\frac{1}{2}+\frac{1}{n}\right),$
so q=2n/(n1).q=2n/(n-1).
Acknowledgements
The first author was supported in part by Gruppo Nazionale per l’Analisi Matematica, by Institute of Mathematics and Informatics, Bulgarian Academy of Sciences and by Top Global University Project, Waseda University. The second author was supported in part by JSPS KAKENHI Grant Number 24H00024.

### v2

@@
 The solution to this problem can be reformulated in terms of the completeness of wave operators.
-For $n=1$, we obtain a complete answer, even for large initial data.
+For $n=1$, we obtain affirmative answer, provided appropriate control on $L^{\infty}$ norm even for large initial data.
 For $n=2$ completeness is established under suitable control of the $L^{\infty}$ norm of the solution $u(t).$
-We prove that the $L^{\infty}$ bound of the solution $u(t)$ matches the decay rate of the linear Schrödinger equation only in the case $n=1.$ For $n\geq 2,$ a similar bound holds only for the $L^{2n/(n-1)}$ norm of $u(t).$
-In the case where no control on the $L^{\infty}$ norm is available, we establish similar asymptotic behavior by replacing the $L^{2}$ norm in (0.1) with smaller Besov-type spaces.
+In the case where no control on the $L^{\infty}$ norm is available, we establish similar asymptotic behavior by replacing the convergence in $L^{2}$ norm in (0.1) with a weaker convergence.
 Keywords: Modified Scattering Nonlinear Schrödinger equation
@@
 The above short overview on existing results shows that most of the results require smallness of the initial data
-Our goal is to solve Q2) without smallness on the initial data at least for $n=1.$
+Our goal is to discuss Q2) without smallness on the initial data for $n=1,2$ or $n\geq 3$ and assuming $L^{\infty}$ control of the solution.
 Remark 1.1.
-Let us mention that in the case $n=1$ the question Q1) is treated in the preprint [25].
+Let us mention that in the case $n=1$ the question Q1) is treated in a preprint [22].
 2. Main results
@@
 $v(0)=u_{0}\in\mathcal{H}^{1,1}.$
-Next, our goal is to prove some $L^{q}$ decay estimates closely connected with the key assumption (2.1) in Theorem 2.1.
+In the case $n=1$ we can control the $L^{\infty}$ norm and this is our next result.
 Theorem 2.3.
-If $n=1$ and $u_{0}\in\mathcal{H}^{1,1},$ then the solution to (1.1) satisfies
-(2.12) $\|u(t)\|_{L^{\infty}}\leq Ct^{-1/2},t\geq 1.$
-For $n\geq 2$ we can control the $L^{q}$ norm of the solution, where $q$ is limited by an index associated with Caffarelli - Kohn - Nirenberg inequality.
+For any $R>0$ there exists $C(R)$ so that for any initial data satisfying
+(2.12) $u_{0}\in\mathcal{H}^{1,1},\ \ \|u_{0}\|_{\mathcal{H}^{1,1}}\leq R$
+the Cauchy problem
+(2.13) $(i\partial_{t}+\frac{1}{2}\Delta_{x})u=u|u|^{2},\ \ t\geq 0;$
+$u(0)=u_{0}\in\mathcal{H}^{1,1},$
+has a unique global solution $u\in C([0,+\infty);\mathcal{H}^{1,1})$ such that
+(2.14) $\|u(t,\cdot)\|_{L^{\infty}}\leq C(R)\ t^{-1/3},\ \ t\geq 1.$
+In the case $n=2$ we can control the $L^{\infty}$ norm under the assumption that initial data are radial.
 Theorem 2.4.
-If $n\geq 2$, $u_{0}\in\mathcal{H}^{1,1},$ and $q\in[2,2n/(n-1)],$ then the solution to (1.1) satisfies
-(2.13) $\|u(t)\|_{L^{q}}\leq Ct^{-n/2+n/q},t\geq 1.$
-We have another estimate provided initial data are radial.
-Theorem 2.5.
-If $n\geq 2$ and $u_{0}\in\mathcal{H}^{1,1}_{rad},$ then the solution to (1.1) satisfies
-(2.14) $\||x|^{n/2}u(t)\|_{L^{\infty}}\leq C,t\geq 1.$
-Again we give a brief idea of the proofs of these $L^{q}$ bounds.
-We use pseudoconformal transformation (2.10). Then the results of Theorems 2.3, 2.4, 2.2 follow from the following.
-Theorem 2.6.
-If $n=1$ and $u_{0}\in\mathcal{H}^{1,1},$ then the solution to (2.11) satisfies
-(2.15) $\|v(t)\|_{L^{\infty}}\leq C,t\in[0,1).$
-Theorem 2.7.
-If $n\geq 2$, $u_{0}\in\mathcal{H}^{1,1}$ and $q\in[2,2n/(n-1)],$ then the solution to (2.11) satisfies
-(2.16) $\|v(t)\|_{L^{q}}\leq C,t\in[0,1).$
-Theorem 2.8.
-If $n\geq 2$ and $u_{0}\in\mathcal{H}^{1,1}_{rad},$ then the solution to (2.11) satisfies
-(2.17) $\||x|^{n/2}v(t)\|_{L^{\infty}}\leq C,t\in[0,1).$
-Then the $L^{q}$ bounds are obtained by using the energy estimate for $v$ (see Lemma 6.2)
-(2.18) $\int_{\mathbb{R}^{n}}|\nabla_{x}v(t,x)|^{2}dx\leq C(1-t)^{-1}$
-and the virial estimate;
-(2.19) $\int_{\mathbb{R}^{n}}|x|^{2}|v(t,x)|^{2}dx\leq C(1-t).$
-A natural interpretation of this estimate is that the solution is spreading out (dispersing) with a very strong rate.
-The crucial estimate (2.19) is obtained via appropriate half - pseudoconformal transformation
-(2.20) $v(t,x)=e^{-\frac{ix^{2}}{4(1-t)}}(1-t)^{-n/4}Y\left(t,\frac{x}{\sqrt{1-t}}\right).$
-Then $Y(t,y)$ satisfies a non - linear Schrödinger equation involving inverted harmonic oscillator.
-We use a combination between virial identity for this equation and energy estimate involving the vector field
-(2.21) $\partial_{y}-i\frac{y}{2}$
-that allows us to obtain
-(2.22) $\int_{\mathbb{R}^{n}}|y|^{2}|Y(t,y)|^{2}dy\leq C$
-and then we arrive at (2.19).
+For $n=2$ any $R>0$ there exists $C(R)$ so that for any initial data satisfying
+(2.15) $u_{0}\in\mathcal{H}^{1,1}_{rad},\ \|u_{0}\|_{\mathcal{H}^{1,1}_{rad}}\leq R$
+the Cauchy problem
+(2.16) $(i\partial_{t}+\frac{1}{2}\Delta_{x})u=u|u|,\ \ t\geq 0;$
+$u(0)=u_{0}\in\mathcal{H}^{1,1}_{rad},$
+has a unique global solution $u\in C([0,+\infty);\mathcal{H}^{1,1}_{rad})$ such that
+(2.17) $|u(t,x)|\leq C(R)\varepsilon^{-1/2}|x|^{-\varepsilon}t^{-(1-\varepsilon)/2}$
+for any $x\in\mathbb{R}^{2}\setminus\{0\},$ any $t>1$ and for any $\varepsilon\in(0,1/2].$
+The idea of the proofs of these $L^{\infty}$ bounds is based on the pseudo - conformal estimate
+(2.18) $t^{-1}\|J(t)u(t)\|_{L^{2}}^{2}\lesssim 1$
+where $t\geq 1$ and
+(2.19) $J(t)=x+it\nabla=U(t)xU(-t)=M(t)it\nabla M(-t).$
 In addition, we turn to the case where no $L^{\infty}$ control is guaranteed.
@@
 Theorem 2.5.
-For any $u_{0}\in\mathcal{F}(H^{1})$ there exists a unique $u_{+}\in L^{2},$ so that
+Let $n=2.$ For any $u_{0}\in\mathcal{F}(H^{1})$ there exists a unique $u_{+}\in L^{2},$ so that
 (2.20) $\lim_{t\to+\infty}\left\|U(-t)\exp\left(i\Phi_{t}(u)\right)u(t)-u_{+}\right\|_{\mathcal{F}(\dot{B}^{-1}_{2,\infty})}=0,$
@@
 Theorem 2.6.
-For any $u_{0}\in\mathcal{F}(H^{1})$ there exists a unique $u_{+}\in L^{2},$ so that
+Let $n\geq 3.$ For any $u_{0}\in\mathcal{F}(H^{1})$ there exists a unique $u_{+}\in L^{2},$ so that
 (2.23) $\lim_{t\to+\infty}\left\|U(-t)\exp\left(i\Phi_{t}(u)\right)u(t)-u_{+}\right\|_{\mathcal{F}(\dot{B}^{-n/2}_{2,\infty})}=0,$
@@
 Passing to $t=\tau/(1-\tau),$ we find $\partial_{\tau}=(1+t)^{2}\partial_{t}$, so
-(6.21) $\frac{1}{4(1+t)}\frac{d}{dt}\|J(1+t)u(t)\|_{L^{2}}^{2}+\frac{d}{dt}\frac{n}{2(n+1)}\|u(t)\|_{L^{2(n+1)/n}}^{2(n+1)/n}=0$
+(6.21) $\frac{(1+t)^{2}}{4}\frac{d}{dt}\left(\frac{1}{1+t}\|J(1+t)u(t)\|_{L^{2}}^{2}\right)$
+$+(1+t)^{2}\frac{d}{dt}\left(\frac{n(1+t)}{2(n+1)}\|u(t)\|_{L^{2(n+1)/n}}^{2(n+1)/n}\right)$
+$+\frac{1}{4}\|J(1+t)u(t)\|_{L^{2}}^{2}=0$
 and
-(6.22) $\frac{d}{dt}\left(\frac{1}{4(1+t)}\|J(1+t)u(t)\|_{L^{2}}^{2}+\frac{n}{2(n+1)}\|u(t)\|_{L^{2(n+1)/n}}^{2(n+1)/n}\right)$
+(6.22) $\frac{d}{dt}\left(\frac{1}{4(1+t)}\|J(1+t)u(t)\|_{L^{2}}^{2}+\frac{n(1+t)}{2(n+1)}\|u(t)\|_{L^{2(n+1)/n}}^{2(n+1)/n}\right)$
 $+\frac{1}{4(1+t)^{2}}\|J(1+t)u(t)\|_{L^{2}}^{2}=0.$
@@
 $+\int_{0}^{\tau}\|\nabla v(\sigma)\|^{2}_{L^{2}}d\sigma=C_{*},$
-$C_{*}=\frac{1}{2}\|\nabla v(0)\|^{2}_{L^{2}}+\frac{2n}{n+1}\|v(0)\|^{2+2/n}_{L^{2+2/n}}.$
+$C_{*}=\|\nabla v(0)\|^{2}_{L^{2}}+\frac{2n}{n+1}\|v(0)\|^{2+2/n}_{L^{2+2/n}}.$
 Proof of Propostition 6.1.
@@
 so we have (2.5).
+8. $L^{\infty}$ bounds
+We assume that $u$ is solution to the Cauchy problem (1.1) with initial data in $\mathcal{H}^{1,1}.$
+We have the following conservation laws.
+  • Charge:
+    $\|u(t)\|_{L^{2}}^{2}=\|u_{0}\|_{L^{2}}^{2}.$
+  • Energy :
+    $\|\nabla u(t)\|_{L^{2}}^{2}+\frac{2n}{n+1}\|u(t)\|_{L^{2+2/n}}^{2+2/n}=\|\nabla u_{0}\|_{L^{2}}^{2}+\frac{2n}{n+1}\|u_{0}\|_{L^{2+2/n}}^{2+2/n}.$
+  • Pseudo - conformal charge:
+    (8.1) $(1+t)^{-1}\|J(1+t)u(t)\|_{L^{2}}^{2}+\frac{2n}{n+1}(1+t)\|u(t)\|_{L^{2+2/n}}^{2+2/n}$
+    $+\int_{0}^{t}(1+s)^{-2}\|J(s+1)u(s)\|_{L^{2}}^{2}ds$
+    $=\|J(1)u(0)\|_{L^{2}}^{2}+\frac{2n}{n+1}\|u(0)\|_{L^{2+2/n}}^{2+2/n},$
+    where
+    (8.2) $J(t)=x+it\nabla=U(t)xU(-t)=M(t)it\nabla M(-t).$
+Proof of Theorem 2.3.
+Now we use the estimates  [^1] We use the following estimates.
+For any $\varphi\in H^{1}(\mathbb{R})$ (8.3) $|\varphi(x)|^{p}=-\int_{x}^{+\infty}\frac{d}{dy}\left(|\varphi(y)|^{p}\right)dy$ $\leq p\int_{-\infty}^{+\infty}|\varphi|^{p-1}|\varphi^{\prime}|dy\leq p\|\varphi\|_{L^{2p-2}}^{p-1}\|\varphi^{\prime}\|_{L^{2}}.\ \$Hence (8.4) $\|\varphi\|_{L^{\infty}}\leq p^{1/p}\|\varphi\|_{L^{2p-2}}^{1-1/p}\|\varphi^{\prime}\|_{L^{2}}^{1/p}.$
+(8.5) $\|u(t)\|_{L^{\infty}}^{3}=\|M(-t)u(t)\|_{L^{\infty}}^{3}\leq 3\|M(-t)u(t)\|^{2}_{L^{4}}\|\nabla M(-t)u\|_{L^{2}}$
+$=3\|u(t)\|^{2}_{L^{4}}\left(t^{-1}\|J(t)u(t)\|_{L^{2}}\right)\leq C(t^{-1/4})^{2}(t^{-1}t^{1/2})=Ct^{-1}$
+for $t\geq 1.$ Hence we have (2.14). ∎
+Proof of Theorem 2.4.
+We take $u_{0}\in\mathcal{H}^{1,1}$ radial.
+Then for any $x\in\mathbb{R}^{2}\setminus\{0\}$, $p\geq 2,$ and $t\geq 1$ we have  [^2] Here for any $\varphi(r)\in H^{1}(\mathbb{R}),$ any $r>0$ and any $p\geq 2$ we use the inequalities (8.6) $|\varphi(r)|^{p}\leq p\left(\int_{r}^{+\infty}|\varphi(p)|^{2p-2}dp\right)^{1/2}\left(\int_{r}^{+\infty}|\varphi^{\prime}(p)|^{2}dp\right)^{1/2}$ $\leq\frac{p}{r}\left(\int_{r}^{+\infty}|\varphi(p)|^{2p-2}pdp\right)^{1/2}\left(\int_{r}^{+\infty}|\varphi^{\prime}(p)|^{2}pdp\right)^{1/2}.$ Hence, for $\varphi\in H^{1}_{rad}(\mathbb{R}^{2})$ and any $x\in\mathbb{R}^{2}\setminus\{0\}$ we have (8.7) $|\varphi(x)|^{p}\leq\frac{p}{|x|}\|\varphi\|_{L^{2p-2}}^{p-1}\|\nabla\varphi\|_{L^{2}}.$ On the other hand, there exists a constant $C\geq 1,$ so that for any $q\in[2,+\infty)$ and any $\varphi\in H^{1}(\mathbb{R}^{2})$ we have (8.8) $\|\varphi\|_{L^{q}}\leq Cq^{1/2}\|\varphi\|_{L^{2}}^{2/q}\|\nabla\varphi\|_{L^{2}}^{1-2/q}.$
+(8.9) $|u(t,x)|^{p}=|M(-t)u(t,x)|^{p}\leq\frac{p}{|x|}\|M(-t)u(t)\|_{L^{2p-2}}^{p-1}\|\nabla M(-t)u(t)\|_{L^{2}}.$
+Using the fact that
+(8.10) $\|M(-t)u(t)\|_{L^{2p-2}}$
+$\leq C(2p-2)^{1/2}\|M(-t)u(t)\|_{L^{2}}^{2/(2p-2)}\|\nabla M(-t)u(t)\|_{L^{2}}^{1-2/(2p-2)},$
+we obtain
+(8.11) $|u(t,x)|^{p}$
+$\leq\frac{p}{|x|}C^{p-1}2^{(p-1)/2}p^{(p-1)/2}\|u_{0}\|_{L^{2}}\|\nabla M(-t)u(t)\|_{L^{2}}^{p-1}$
+$\leq\frac{p}{|x|}C^{p-1}2^{(p-1)/2}p^{(p-1)/2}\|u_{0}\|_{L^{2}}\left(t^{-1}\|J(t)u(t)\|_{L^{2}}\right)^{p-1}$
+$\leq\frac{p}{|x|t^{p-1}}C^{p}(2p)^{p/2}\|u_{0}\|_{L^{2}}\|J(t)u(t)\|_{L^{2}}^{p-1}.$
+for any $t\geq 1.$ In this way, we find
+(8.12) $|u(t,x)|\leq\left(\frac{p}{|x|}\|u_{0}\|_{L^{2}}\right)^{1/p}C(2p)^{1/2}t^{-(p-1)/(2p)}.$
+Substitution $\varepsilon=1/p$ implies (2.17).
 9. Weak convergence of modified wave operators
@@
 (9.4) $\mathcal{F}M(t)U(-t)(gu(t))=\tilde{g}\mathcal{F}M(t)U(-t)u(t),\ \ \ \tilde{g}(x)=g(xt).$
-Then we have (4.10) and we can refer to Corolary 4.1 that implies
+Then we have (4.10) and we can refer to Corollary 4.1 that implies
 (9.5) $i\frac{d}{dt}v(t)=\mathcal{F}^{-1}\exp\left(i\int_{0}^{t}|u(\tau,\tau x)|d\tau\right)\frac{(-\Delta)}{2t^{2}}\mathcal{F}M(t)U(-t)u(t).$
@@
 (10.27) $\tilde{\Phi}_{t}(x)=\int_{1}^{t}\left[(1+|u(\tau,\tau x)|^{2})^{1/n}-1\right]d\tau.$
-Then we have (4.10) and we can refer to Corolary 4.1 that implies
+Then we have (4.10) and we can refer to Corollary 4.1 that implies
 (10.28) $i\frac{d}{dt}v(t)=\mathcal{F}^{-1}\exp\left(i\tilde{\Phi}_{t}(x)\right)\frac{(-\Delta)}{2t^{2}}\mathcal{F}M(t)U(-t)u(t).$
@@
 This completes the proof.
-10. Virial bound after pseudoconformal transformation
-We continue assuming $v$ is a solution to (6.6), replacing $(\tau,\xi)$ by $(t,x)$, we have
-(10.1) $(i\partial_{t}+\frac{1}{2}\Delta_{x})v=\frac{v(t)|v(t)|^{2/n}}{1-t},\ \ \ t\in[0,1);$
-$v(0)=v_{0}\in\mathcal{H}^{1,1}.$
-Lemma 10.1.
-For any $n\geq 1$ and any $R>0$ there exists a constant $C=C(R)>0$ so that if $v$ solves the problem (10.1) with initial data satisfying
-(10.2) $\|v_{0}\|_{H^{1}}+\||x|v_{0}\|_{L^{2}}\leq R,$
-then $v$ satisfies the inequality
-(10.3) $\int_{\mathbb{R}^{n}}|x|^{2}|v(t,x)|^{2}dx\leq C(R).$
-Proof.
-It is sufficient to prove (10.3) assuming $v_{0}\in\mathcal{H}^{N,1}$ with $N>n/2+2.$ Then $v\in C([0,1);\mathcal{H}^{N,1}).$ We have the following continuity equation, obtained by multiplying (10.1) by $\overline{v}$ and taking the imaginary part
-(10.4) $\partial_{t}|v(t,x)|^{2}+\nabla_{x}\left[\mathrm{Im}\left(\nabla_{x}v(t,x)\overline{v(t,x)}\right)\right]=0.$
-Now we introduce the following momentum associated with $v$
-(10.5) $M_{v}(t)=\int_{\mathbb{R}^{n}}|x|^{2}|v(t,x)|^{2}dx.$
-Multiplying the equation (10.4) by the weight $|x|^{2}$ and integrating by parts in $x$ we find
-(10.6) $M_{v}(t)\leq M_{v}(0)+2\int_{0}^{t}\sqrt{M_{v}(\tau)}\|\nabla_{x}v(\tau)\|_{L^{2}}d\tau.$
-From this estimate we can get
-(10.7) $M_{v}(t)\leq M_{v}(0)$
-$+2\left(\sup_{0\leq\tau\leq t}M_{v}(\tau)\right)^{1/2}\sup_{0\leq\tau\leq t}(1-\tau)^{1/2}\|\nabla_{x}v(\tau)\|_{L^{2}}\int_{0}^{t}(1-\tau)^{-1/2}d\tau$
-$\leq M_{v}(0)+\frac{1}{2}\left(\sup_{0\leq\tau\leq t}M_{v}(\tau)\right)+2\left(\sup_{0\leq\tau\leq t}(1-\tau)^{1/2}\|\nabla_{x}v(\tau)\|_{L^{2}}\right)^{2}.$
-Recall the energy identity (6.25) that implies
-(10.8) $\sup_{0\leq\tau\leq t}(1-\tau)^{1/2}\|\nabla_{x}v(\tau)\|_{L^{2}}\leq C_{1}(R)$
-and we deduce
-(10.9) $\sup_{0\leq\tau\leq t}M_{v}(\tau)\leq M_{v}(0)+\frac{1}{2}\left(\sup_{0\leq\tau\leq t}M_{v}(\tau)\right)+2C_{1}(R)^{2}$
-so
-(10.10) $M_{v}(t)\leq 2M_{v}(0)+4C_{1}(R)^{2}.$
-This completes the proof.
-11. Transformation of NLS to nonlinear problem for inverted harmonic oscillator
-In this section, we continue to study the Cauchy problem (10.1), so that we are looking for a transformation of the NLS in (10.1) to a new nonlinear evolution problem associated with appropriate Hamiltonian such that the uniform bounds
-(11.1) $\|v(t)\|_{L^{2}}=const,$
-$(1-t)^{1/2}\|\nabla_{x}v(t)\|_{L^{2}}\lesssim 1$
-can be interpreted as mass and kinetic energy for the new Hamiltonian. We make the following substitution $(t,x,v)\mapsto(s,y,Y)$ defined by
-(11.2) $s=t,\ \ y=\frac{x}{(1-t)^{1/2}},$
-$Y(s,y)=e^{iy^{2}/4}(1-s)^{n/4}v(s,y(1-s)^{1/2})$
-with inverse substitution given by
-(11.3) $t=s,\ x=y(1-s)^{1/2},$
-$v(t,x)=e^{-\frac{ix^{2}}{4(1-t)}}(1-t)^{-n/4}Y\left(t,\frac{x}{\sqrt{1-t}}\right).$
-Lemma 11.1.
-The transformation (11.2), namely
-(11.4) $Y(s,y)=e^{iy^{2}/4}(1-s)^{n/4}v(s,y(1-s)^{1/2})$
-transforms the equation
-(11.5) $\left(i\partial_{t}+\frac{\Delta_{x}}{2}\right)v(t,x)=\frac{v(t,x)|v(t,x)|^{2/n}}{1-t},$
-into
-(11.6) $i(1-s)\partial_{s}Y(s,y)-\mathcal{H}_{1/2}Y(s,y)=\frac{Y(s,y)|Y(s,y)|^{2/n}}{(1-s)^{1/2}},$
-where $\mathcal{H}_{\omega}$ is inverted harmonic oscillator Hamiltonian
-(11.7) $\mathcal{H}_{\omega}=-\frac{1}{2}\Delta_{y}-\frac{\omega^{2}}{2}y^{2}.$
-Proof.
-A direct substitution
-(11.8) $v(t,x)=e^{-\frac{ix^{2}}{4(1-t)}}(1-t)^{-n/4}\ Y\left(t,\frac{x}{\sqrt{1-t}}\right)$
-gives
-(11.9) $(1-t)^{1+n/4}e^{\frac{ix^{2}}{4(1-t)}}\left(i\partial_{t}+\frac{\Delta_{x}}{2}\right)v(t,x)$
-$=\left[\left(i(1-t)\partial_{t}Y+\frac{\Delta_{y}}{2}Y\right)\left(t,\frac{x}{\sqrt{1-t}}\right)+\frac{x^{2}}{8(1-t)}Y\left(t,\frac{x}{\sqrt{1-t}}\right)\right].$
-On the other hand, we have
-(11.10) $e^{\frac{ix^{2}}{4(1-t)}}(1-t)^{1+n/4}\frac{v(t,x)|v(t,x)|^{2/n}}{1-t}$
-$=(1-t)^{-1/2}Y\left(t,\frac{x}{\sqrt{1-t}}\right)\left|Y\left(t,\frac{x}{\sqrt{1-t}}\right)\right|^{2/n}.$
-Hence,
-(11.11) $e^{\frac{ix^{2}}{4(1-t)}}(1-t)^{1+n/4}\left[\left(i\partial_{t}+\frac{\Delta_{x}}{2}\right)v(t,x)-\frac{v(t,x)|v(t,x)|^{2/n}}{1-t}\right]$
-$=\left[\left(i(1-t)\partial_{t}Y+\frac{\Delta_{y}}{2}Y\right)\left(t,\frac{x}{\sqrt{1-t}}\right)+\frac{x^{2}}{8(1-t)}Y\left(t,\frac{x}{\sqrt{1-t}}\right)\right]$
-$-(1-t)^{-1/2}\ Y\left(t,\frac{x}{\sqrt{1-t}}\right)\left|Y\left(t,\frac{x}{\sqrt{1-t}}\right)\right|^{2/n}.$
-Lemma 11.2.
-Let $v(t,x)$ be a solution to the Cauchy problem (10.1) and $Y(s,y)$ is defined by (11.2). Then we have the following properties.
-  1. i)
-    We have the relations
-    (11.12) $\|Y(t)\|^{2}_{L^{2}}=\|v(t)\|_{L^{2}}^{2},$
-    $\|Y(t)\|^{2(n+1)/n}_{L^{2(n+1)/n}}=(1-t)^{1/2}\|v(t)\|^{2(n+1)/n}_{L^{2(n+1)/n}}.$
-  2. ii)
-    We have the uniform bounds
-    (11.13) $\left\|\left(\nabla_{y}-\frac{iy}{2}\right)Y(t)\right\|^{2}_{L^{2}}\lesssim 1,$
-    $(1-t)\||y|Y(t)\|^{2}_{L^{2}}=\||x|v(t)\|^{2}_{L^{2}}\lesssim 1.$
-Proof.
-We use (11.4) and can write
-(11.14) $\|Y(t)\|^{2}_{L^{2}}=\int_{\mathbb{R}^{n}}|v(t,y(1-t)^{1/2})|^{2}(1-t)^{n/2}dy=\|v(t)\|^{2}_{L^{2}},$
-and
-(11.15) $\|Y(t)\|^{q}_{L^{q}}=\int_{\mathbb{R}^{n}}|v(t,y(1-t)^{1/2})|^{2}(1-t)^{qn/4}dy$
-$=(1-t)^{qn/4-n/2}\|v(t)\|^{q}_{L^{q}},$
-with $q=2(n+1)/n$ we find
-(11.16) $\|Y(t)\|^{2(n+1)/n}_{L^{2(n+1)/n}}=(1-t)^{2(n+1)/4-n/2}\|v(t)\|^{2(n+1)/n}_{L^{2(n+1)/n}}$
-$=(1-t)^{1/2}\|v(t)\|^{2(n+1)/n}_{L^{2(n+1)/n}}.$
-This proves (11.12).
-Now we use the definition (11.4) of $Y$ and find
-(11.17) $\left(\nabla_{y}-\frac{iy}{2}\right)Y(s)$
-$=\left(\nabla_{y}-\frac{iy}{2}\right)\left(e^{iy^{2}/4}(1-s)^{n/4}v(s,y(1-s)^{1/2})\right)$
-$=e^{iy^{2}/4}(1-s)^{n/4+1/2}\nabla v(s,y(1-s)^{1/2}).$
-Hence,
-(11.18) $\left\|\left(\nabla_{y}-\frac{iy}{2}\right)Y(s)\right\|^{2}_{L^{2}}$
-$=\int_{\mathbb{R}^{n}}(1-s)^{n/2+1}|\nabla v(s,y(1-s)^{1/2})|^{2}dy$
-$=(1-s)\|\nabla v(s)\|_{L^{2}}^{2}.$
-Now we apply the energy estimate
-(11.19) $(1-s)\|\nabla v(s)\|_{L^{2}}^{2}\lesssim 1$
-following from the energy identity (6.25) in Lemma 6.2 and deduce the corresponding inequality in (11.13).
-Finally, we have
-(11.20) $(1-t)\||y|Y(t)\|^{2}_{L^{2}}$
-$=\int_{\mathbb{R}^{n}}(1-t)|y|^{2}|v(t,y(1-t)^{1/2})|^{2}(1-t)^{n/2}dy$
-$=\||x|v(t)\|^{2}_{L^{2}}.$
-Now we can use Lemma 10.1 and see that
-(11.21) $(1-t)\||y|Y(t)\|^{2}_{L^{2}}\lesssim 1.$
-This completes the proof of (11.13) and the Lemma. ∎
-12. Virial estimates for inverted harmonic oscillator
-12.1. On linear inverted harmonic oscillator
-We have the following information for the Hamiltonian known also a inverted harmonic oscillator
-(12.1) $\mathcal{H}_{\omega}=-\frac{1}{2}\Delta_{y}-\frac{\omega^{2}}{2}y^{2},\ \omega>0:$
-  1. i)
-    The operator $\mathcal{H}_{\omega}$ with domain
-    (12.2) $\Sigma_{2}=\{f\in H^{2},|y|^{2}f\in L^{2}\}$
-    is self - adjoint with absolutely continuous spectrum
-    (12.3) $\sigma(\mathcal{H}_{\omega})=\sigma_{ac}(\mathcal{H}_{\omega})=\mathbb{R}.$
-  2. ii)
-    The operator has no resonances. Namely, the result in [23] guarantees that any solution to
-    (12.4) $(\mathcal{H}_{\omega}-\lambda)U=0$
-    with
-    $U\in\langle x\rangle^{s}L^{2},\ \ s<\frac{1}{2}$
-    has to be the trivial one $U=0.$
-  3. iii)
-    We have limiting absorption principle and $L^{2}$ scattering for appropriate perturbations of $\mathcal{H}_{\omega}$ (see [3], [23]).
-12.2. Virial bound for inverted harmonic oscillator
-We assume that $Y(s,y),s\in[0,1),y\in\mathbb{R}^{n}$ is a solution to the Cauchy problem
-(12.5) $i(1-s)\partial_{s}Y(s,y)-\mathcal{H}_{1/2}Y(s,y)=\frac{Y(s,y)|Y(s,y)|^{2/n}}{(1-s)^{1/2}},$
-$Y(0)=Y_{0}\in\mathcal{H}^{1,1}.$
-Lemma 12.1.
-(Morawetz identity: first step) We assume $Y$ is a solution to (12.5). Then we have
-(12.6) $\frac{(1-t)}{2}M_{Y}^{\prime}(t)=\frac{1}{2}\mathrm{Im}\left(\left.LY(t)\right|\overline{Y(t)}\right),$
-where
-(12.7) $\left(\left.f\right|g\right)=\int_{\mathbb{R}^{n}}f(y)\overline{g(y)}dy,$
-(12.8) $M_{Y}(t)=\int_{\mathbb{R}^{n}}y^{2}|Y(t,y)|^{2}dy$
-and
-(12.9) $L=y\cdot\nabla_{y}+\nabla_{y}\cdot y$
-is the dilation generator.
-Proof.
-Again as in the proof of Lemma 10.1 it is sufficient to prove (12.6) assuming $Y_{0}\in\mathcal{H}^{N,1}$ with $N>n/2+2.$ Then $Y\in C([0,1);\mathcal{H}^{N,1}).$
-We have the following continuity equation obtained by multiplying (11.6) by $\overline{Y}$ and taking the imaginary part
-(12.10) $\frac{(1-t)}{2}\partial_{t}|Y(t,y)|^{2}+\frac{1}{2}\nabla_{y}\left[\mathrm{Im}\left(\nabla_{y}Y(t,y)\overline{Y(t,y)}\right)\right]=0.$
-Multiplying by $y^{2}$ and integrating in $y$, we find
-(12.11) $0=\frac{(1-t)}{2}\partial_{t}\left(\int_{\mathbb{R}^{n}}y^{2}|Y(t,y)|^{2}dy\right)-\left(\int_{\mathbb{R}^{n}}\mathrm{Im}\left(y\cdot\nabla_{y}Y(t,y)\overline{Y(t,y)}\right)dy\right)$
-$=\frac{(1-t)}{2}\partial_{t}\left(\int_{\mathbb{R}^{n}}y^{2}|Y(t,y)|^{2}dy\right)-\frac{1}{2}\int_{\mathbb{R}^{n}}\mathrm{Im}\left(y\cdot\nabla_{y}Y(t,y)\overline{Y(t,y)}\right)dy$
-$+\frac{1}{2}\int_{\mathbb{R}^{n}}\mathrm{Im}\left(Y(t,y)\overline{y\cdot\nabla_{y}Y(t,y)}\right)dy$
-$=\frac{(1-t)}{2}\partial_{t}\left(\int_{\mathbb{R}^{n}}y^{2}|Y(t,y)|^{2}dy\right)-\frac{1}{2}\mathrm{Im}\left(\left.\left(y\cdot\nabla_{y}+\nabla_{y}\cdot y\right)Y(t)\right|Y(t)\right).$
-Lemma 12.2.
-There exists a constant $C_{0}\in\mathbb{R}$ so that we have the asymptotic expansion
-(12.12) $M_{Y}(t)=\frac{C_{0}}{1-t}+rem(t),$
-$|rem(t)|\leq C(1-t)^{-1/2},$
-$|rem^{\prime}(t)|\leq C(1-t)^{-3/2}.$
-Proof.
-We set
-(12.13) $Z(t)=(1-t)M_{Y}(t)$
-and applying Lemma 12.1, we have
-(12.14) $\frac{1}{2}Z^{\prime}(t)=\frac{1-t}{2}M_{Y}^{\prime}(t)-\frac{1}{2}M_{Y}(t),$
-$=\frac{1}{2}\mathrm{Im}\left(\int_{\mathbb{R}^{n}}\left[(y\cdot\nabla_{y}+\nabla_{y}\cdot y)Y(t,y)\right]\,\overline{Y(t,y)}dy\right)$
-$-\frac{1}{2}\mathrm{Im}\left(\int_{\mathbb{R}^{n}}\left[(iy\cdot y+iy\cdot y)Y(t,y)\right]\ \overline{Y(t,y)}dy\right)$
-$=\frac{1}{2}\mathrm{Im}\left(\int_{\mathbb{R}^{n}}\left[(y\cdot(\nabla_{y}-iy)+(\nabla_{y}-iy)\cdot y)Y(t,y)\right]\ \overline{Y(t,y)}dy\right).$
-Therefore, we use the estimates of Lemma 11.2 and find
-(12.15) $|Z^{\prime}(t)|\leq C\sqrt{M_{Y}(t)}\leq C(1-t)^{-1/2}.$
-We know that $Z(t),t\in[0,1)$ is bounded. Therefore, there exists $t_{k}\nearrow 1$ so that
-(12.16) $\lim_{k\to\infty}Z(t_{k})=C_{0}.$
-From this and the estimate of $Z^{\prime}$ it is easy to see that $\lim_{t\nearrow 1}Z(t)$ exists, we denote this limit by $C_{0}.$ Then we can write
-(12.17) $Z(t)=C_{0}-\int_{t}^{1}Z^{\prime}(\tau)d\tau,$
-and then from
-(12.18) $rem(t)=-\frac{1}{1-t}\int_{t}^{1}Z^{\prime}(\tau)d\tau,$
-we get
-(12.19) $|rem(t)|\leq C(1-t)^{-1/2}.$
-From (12.18) we can estimate the derivative of $rem(t)$ by using (12.15), so we can write
-(12.20) $|rem^{\prime}(t)|\leq\left|\frac{1}{(1-t)^{2}}\int_{t}^{1}Z^{\prime}(\tau)d\tau\right|$
-$+\left|\frac{1}{1-t}Z^{\prime}(t)d\tau\right|\lesssim(1-t)^{-3/2}.$
-Lemma 12.3.
-(Morawetz identity: second step) If $Y$ is a solution to
-(12.21) $i(1-t)\partial_{t}Y(t,y)-\mathcal{H}_{\omega}Y(t,y)=\frac{Y(t,y)|Y(t,y)|^{2/n}}{(1-t)^{1/2}},$
-where $\mathcal{H}_{\omega},\omega>0$ is inverted harmonic oscillator Hamiltonian
-(12.22) $\mathcal{H}_{\omega}=-\frac{1}{2}\Delta_{y}-\frac{\omega^{2}}{2}y^{2},$
-then we have
-(12.23) $\frac{1}{4}\left((1-t)\frac{d}{dt}\right)^{2}M_{Y}(t)=\|\nabla_{y}Y(t)\|_{L^{2}}^{2}+\omega^{2}M_{Y}(t)$
-$+\frac{n}{(n+1)\sqrt{1-t}}\|Y(t)\|_{L^{2+2/n}}^{2+2/n}.$
-Proof.
-We use the dilation generator $L$ defined in (12.9) which is commonly used in virial and Morawetz identities. Then, we have the commutator relation
-(12.24) $\left[\mathcal{H}_{\omega},L\right]=-\left[\frac{1}{4}\Delta+\frac{1}{4}\omega^{2}|y|^{2},y\cdot\nabla_{y}+\nabla_{y}\cdot y\right]=-\Delta+\omega^{2}y^{2}.$
-Here we use the commutator relations
-(12.25) $\left[\Delta,(y\cdot\nabla_{y}+\nabla_{y}\cdot y)\right]=4\Delta,$
-$\left[y^{2},(y\cdot\nabla_{y}+\nabla_{y}\cdot y)\right]=-4y^{2}.$
-Using equation (11.6), we apply the operator $L=y\cdot\nabla_{y}+\nabla_{y}\cdot y$ and find
-(12.26) $i(1-t)\partial_{t}LY(t,y)-L\mathcal{H}_{\omega}Y(t,y)=L\left(\frac{Y(t,y)|Y(t,y)|^{2/n}}{(1-t)^{1/2}}\right)$
-We take the real part of the inner product in $L^{2}$ of the equation in (12.26) and $Y$ and find
-(12.27) $-\frac{1-t}{2}\frac{d}{dt}\mathrm{Im}\left(\left.LY(t)\right|Y(t)\right)-\mathrm{Re}\left(\left.LH_{\omega}Y(t)\right|Y(t)\right)$
-$=\mathrm{Re}\left(\left.L\left(\frac{Y(t)|Y(t)|^{2/n}}{\sqrt{1-t}}\right)\right|Y(t)\right),$
-since
-(12.28) $\frac{d}{dt}\mathrm{Im}\left(\left.LY(t)\right|Y(t)\right)=\mathrm{Im}\left(\left.L\partial_{t}Y(t)\right|Y(t)\right)$
-$+\mathrm{Im}\left(\left.LY(t)\right|\partial_{t}Y(t)\right)=2\mathrm{Im}\left(\left.L\partial_{t}Y(t)\right|Y(t)\right).$
-The second term in (12.27) can be transformed by using (12.24) as follows
-(12.29) $-\mathrm{Re}\left(\left.L\mathcal{H}_{\omega}Y(t)\right|Y(t)\right)$
-$=-\frac{1}{2}\left(\left.L\mathcal{H}_{\omega}Y(t)\right|Y(t)\right)-\frac{1}{2}\overline{\left(\left.L\mathcal{H}_{\omega}Y(t)\right|Y(t)\right)}$
-$=-\frac{1}{2}\left(\left.(L\mathcal{H}_{\omega}-\mathcal{H}_{\omega}L)Y(t)\right|Y(t)\right)$
-$=-\frac{1}{2}\left(\left.[L,\mathcal{H}_{\omega}]Y(t)\right|Y(t)\right)$
-$=-\frac{1}{4}\left(\left.[(\Delta+\omega^{2}y^{2}),(y\cdot\nabla_{y}+\nabla_{y}\cdot y)]Y(t))\right|Y(t)\right)$
-$=\left(\left.(-\Delta+\omega^{2}y^{2})Y(t)\right|Y(t)\right).$
-The last term in (12.27) can be transformed by using the following relation
-(12.30) $\mathrm{Re}\left(\left(Y(t,y)|Y(t,y)|^{2/n}\right)\nabla_{y}\overline{Y(t,y)}\right)=\frac{n}{2(n+1)}\nabla_{y}|Y(t,y)|^{2+2/n}$
-so we have
-(12.31) $\mathrm{Re}\left(\left.L\left(Y(t,y)|Y(t,y)|^{2/n}\right)\right|Y(t)\right)$
-$=2\mathrm{Re}\left(\int_{\mathbb{R}^{n}}\left[\nabla_{y}\cdot y\left(Y(t,y)|Y(t,y)|^{2/n}\right)\right]\overline{Y(t,y)}dy\right)-n\|Y(t)\|_{L^{2+2/n}}^{2+2/n}$
-$=-2\left(\int_{\mathbb{R}^{n}}y\cdot\mathrm{Re}\left(Y(t,y)|Y(t,y)|^{2/n}\right)\nabla_{y}\overline{Y(t,y)}dy\right)-n\|Y(t)\|_{L^{2+2/n}}^{2+2/n}$
-$=-\frac{n}{n+1}\int_{\mathbb{R}^{n}}y\cdot\nabla_{y}(|Y(t,y)|^{2+2/n})dy-n\|Y(t)\|_{L^{2+2/n}}^{2+2/n}$
-$=\left(\frac{n^{2}}{n+1}-n\right)\|Y(t)\|_{L^{2+2/n}}^{2+2/n}=-\frac{n}{n+1}\|Y(t)\|_{L^{2+2/n}}^{2+2/n}.$
-Now (12.27) leads to
-(12.32) $\frac{1-t}{2}\frac{d}{dt}\mathrm{Im}\left(\int_{\mathbb{R}^{n}}LY(t,y)\overline{Y(t,y)}dy\right)$
-$=\int_{\mathbb{R}^{n}}((-\Delta+\omega^{2}y^{2})Y(t,y))\overline{Y(t,y)}dy+\frac{n}{(n+1)\sqrt{1-t}}\|Y(t)\|_{L^{2+2/n}}^{2+2/n}$
-$=\|\nabla_{y}Y(t)\|_{L^{2}}^{2}+\omega^{2}M_{Y}(t)+\frac{n}{(n+1)\sqrt{1-t}}\|Y(t)\|_{L^{2+2/n}}^{2+2/n}.$
-Turning back to (12.6), we obtain (12.23).
-Lemma 12.4.
-The constant $C_{0}$ from Lemma 12.2 is identically zero.
-Proof.
-We estimate the right side of (12.23), by using the estimates of Lemma 11.2.
-(12.33) $\|\nabla_{y}Y(t)\|_{L^{2}}^{2}+\omega^{2}M_{Y}(t)+\frac{n}{(n+1)\sqrt{1-t}}\|Y(t)\|_{L^{2+2/n}}^{2+2/n}$
-$=2\omega^{2}M_{Y}(t)+O(\sqrt{M_{Y}(t)})+O(1).$
-We further use the quantity $Z(t)$ introduced in (12.13). We use the simple relation
-(12.34) $(1-t)\frac{d}{dt}M_{Y}(t)=Z^{\prime}(t)+M_{Y}(t)$
-and find for the left side of (12.23)
-(12.35) $\frac{1}{4}\left((1-t)\frac{d}{dt}\right)^{2}M_{Y}(t)=\frac{1}{4}\left((1-t)\frac{d}{dt}\right)(Z^{\prime}(t)+M_{Y}(t))$
-$=\frac{1}{4}(1-t)Z^{\prime\prime}(t)+\frac{1}{4}Z^{\prime}(t)+\frac{1}{4}M_{Y}(t).$
-From (12.23) we find with $\omega=1/2$
-(12.36) $\frac{1}{4}(1-t)Z^{\prime\prime}(t)=\frac{1}{4}M_{Y}(t)+O((1-t)^{-1/2}),$
-where we have used estimate (12.15). Now we are in position to use the asymptotic expansion (12.12) for $M_{Y}(t)$ and deduce
-(12.37) $Z^{\prime\prime}(t)=\frac{C_{0}}{(1-t)^{2}}+O((1-t)^{-3/2}).$
-Integrating, we get
-(12.38) $Z^{\prime}(t)=Z^{\prime}(0)+\int_{0}^{t}\left(\frac{C_{0}}{(1-\tau)^{2}}+O((1-\tau)^{-3/2})\right)d\tau$
-$=\frac{C_{0}}{1-t}+O((1-t)^{-1/2}).$
-This representation and (12.15) imply $C_{0}=0.$
-Lemma 12.5.
-We have
-(12.39) $M_{Y}(t)\lesssim 1,\ \forall t\in[0,1).$
-Proof.
-We use the inequality (12.15) so we can write
-(12.40) $|Z^{\prime}(t)|\leq C\sqrt{Z(t)}(1-t)^{-1/2}.$
-Integrating with $0<t_{1}<t_{2}<1$, we get
-(12.41) $\sqrt{Z(t_{2})}-\sqrt{Z(t_{1})}\geq-C\left(\sqrt{1-t_{1}}-\sqrt{1-t_{2}}\right).$
-We know that $Z(t_{2})\to 0$ when $t_{2}\to 1,$ so letting $t_{2}\to 0$ we get
-(12.42) $C\sqrt{1-t_{1}}\geq\sqrt{Z(t_{1})}$
-and we arrive at
-(12.43) $Z(t_{1})\leq C(1-t_{1}).$
-This completes the proof.
-From the boundedness of $M_{y}(t)$ and relation (11.13) in Lemma 11.2 we obtain the estimate
-Lemma 12.6.
-If $v$ solves the Cauchy problem (10.1), i.e.
-(12.44) $\left(i\partial_{t}+\frac{\Delta_{x}}{2}\right)v(t,x)=\frac{v(t,x)|v(t,x)|^{2/n}}{1-t},\ \ t\in(0,1),x\in\mathbb{R}^{n}$
-$v(0,x)=v_{0}(x)\in\mathcal{H}^{1,1},$
-then we have the uniform bound
-(12.45) $\int_{\mathbb{R}^{n}}|x|^{2}|v(t,x)|^{2}dx\leq C(1-t).$
-12.3. Proofs of Theorems 2.6, 2.7 and 2.8.
-Proof of Theorem 2.6 .
-The proof for $n=1$ is based on the following simple inequality
-(12.46) $|f(x)|^{2}\leq\frac{2}{x}\left(\int_{x}^{\infty}|f^{\prime}(y)|^{2}dy\right)^{1/2}\left(\int_{x}^{\infty}|f(y)|^{2}y^{2}dy\right)^{1/2},$
-where $x>0.$
-The key point is the following uniform bound (see Lemma 12.6)
-(12.47) $\int_{\mathbb{R}^{n}}|x|^{2}|v(t,x)|^{2}dx\leq C(1-t).$
-The energy estimate (see Lemma 6.2)
-(12.48) $\int_{\mathbb{R}^{n}}|\nabla_{x}v(t,x)|^{2}dx\leq C(1-t)^{-1}.$
-combined with (12.46) and (12.47) imply (2.15) for $x>1.$ Similarly we proceed for x<1.x<-1. Finally, for $|x|\leq 1$ we use the translation invariance (in $x$) of the equation
-$(i\partial_{t}+\frac{1}{2}\Delta_{x})v=\frac{v|v|^{2/n}}{1-t}$
-and complete the proof of Theorem 2.6.
-Proof of Theorem 2.7 .
-For Theorem 2.7 we use Caffarelli - Kohn - Nirenberg inequality ( see (13.1) that is obtained as partial case of more general inequality obtained in [4]).
-(12.49) $\left\||x|^{1/2}u\right\|^{2}_{L^{q}}\lesssim\||\nabla u\|_{L^{2}}\||x|u\|_{L^{2}}.$
-Using the translation invariance of the Schrödinger equation for $v$, we can obtain the following variant of (12.47), namely
-(12.50) $\int_{\mathbb{R}^{n}}|x-x_{0}|^{2}|v(t,x)|^{2}dx\lesssim(1-t)$
-provided $|x_{0}|=1.$ In this way we get
-(12.51) $\left\||x-x_{0}|^{1/2}u\right\|^{2}_{L^{q}}\lesssim\||\nabla u\|_{L^{2}}\||x-x_{0}||u\|_{L^{2}}\lesssim 1.$
-From
-(12.52) $|x-x_{0}|^{1/2}+|x|^{1/2}\gtrsim 1$
-we complete the proof. ∎
-Proof of Theorem 2.8 .
-For Theorem 2.8 we use the following variant of (12.46).
-(12.53) $r^{n-1}|f(r)|^{2}$ $\leq\frac{2}{r}\left(\int_{r}^{\infty}|f^{\prime}(\tau)|^{2}\tau^{n-1}d\tau\right)^{1/2}\left(\int_{r}^{\infty}|f(\tau)|^{2}\tau^{2}\tau^{n-1}d\tau\right)^{1/2}$
-$+\frac{n-1}{r}\int_{r}^{\infty}|f(\tau)|^{2}\tau^{n-1}d\tau.$
-13. Caffarelli - Kohn - Nirenberg inequality
-We recall the classical Caffarelli - Kohn - Nirenberg inequality (see [4]) involving parameters
-$p_{1},p_{2}\geqslant 1,\quad q>0,\quad 0\leqslant a\leqslant 1$
-$\frac{1}{p_{1}}+\frac{\alpha}{n},\quad\frac{1}{p_{2}}+\frac{\beta}{n},\quad\frac{1}{q}+\frac{\gamma}{n}>0,$
-where
-$\gamma=a\sigma+(1-a)\beta.$
-If we require $a\in(0,1)$, then the main result in [4] implies the following.
-Theorem 13.1.
-If
-$0\leqslant\alpha-\sigma$
-and
-$\frac{1}{q}+\frac{\gamma}{n}=a\left(\frac{1}{p_{1}}+\frac{\alpha-1}{n}\right)+(1-a)\left(\frac{1}{p_{2}}+\frac{\beta}{n}\right),$
-then there exists a positive constant $C$ such that the following inequality holds for all $u\in C_{0}^{\infty}\left(\mathbb{R}^{n}\right)$
-$\left\||x|^{\gamma}u\right\|_{L^{q}}\leqslant C\||x|^{\alpha}\nabla u\|_{L^{p_{1}}}^{a}\||x|^{\beta}u\|_{L^{p_{2}}}^{1-a}$
-In the special case $a=1/2,$ $p_{1}=p_{2}=2$, $\alpha=0,$ $\beta=1$ we take $\sigma=\alpha=0$ and deduce $\gamma=1/2$ and
-(13.1) $\left\||x|^{1/2}u\right\|^{2}_{L^{q}}\lesssim\|\nabla u\|_{L^{2}}\||x|u\|_{L^{2}}$
-provided
-(13.2) $\frac{1}{q}+\frac{1}{2n}=\frac{1}{2}\left(\frac{1}{2}-\frac{1}{n}\right)+\frac{1}{2}\left(\frac{1}{2}+\frac{1}{n}\right),$
-so q=2n/(n1).q=2n/(n-1).
 Acknowledgements
+The authors are grateful to J.Murphy, H.Mizutani and H.Miyazaki for the discussions and for pointing out a mistake in the previous version of the manuscript.
 The first author was supported in part by Gruppo Nazionale per l’Analisi Matematica, by Institute of Mathematics and Informatics, Bulgarian Academy of Sciences and by Top Global University Project, Waseda University.
-The second author was supported in part by JSPS KAKENHI Grant Number 24H00024.
+The second author was supported by JSPS KAKENHI Grant Number 24H00024.